### v2

# Visual Instruction Pretraining for Domain-Specific Foundation Models

Abstract
Modern computer vision is converging on a closed loop in which perception, reasoning and generation mutually reinforce each other. However, this loop remains incomplete: the top-down influence of high-level reasoning on the foundational learning of low-level perceptual features is not yet underexplored. This paper addresses this gap by proposing a new paradigm for pretraining foundation models in downstream domains. We introduce Visual insTruction Pretraining (ViTP), a novel approach that directly leverages reasoning to enhance perception. ViTP embeds a Vision Transformer (ViT) backbone within a Vision-Language Model and pretrains it end-to-end using a rich corpus of visual instruction data curated from target downstream domains. ViTP is powered by our proposed Visual Robustness Learning (VRL), which compels the ViT to learn robust and domain-relevant features from a sparse set of visual tokens. Extensive experiments on 16 challenging remote sensing and medical imaging benchmarks demonstrate that ViTP establishes new state-of-the-art performance across a diverse range of downstream tasks. The code is available at https://github.com/zcablii/ViTP.
1 Introduction
Figure 1: The synergistic relationship between perception, generation, and reasoning in modern CV. Our proposed ViTP forges a novel link from high-level reasoning to low-level perception, a previously underexplored connection.
Recent years have witnessed a leap forward in computer vision (CV), largely catalyzed by the Transformer architecture [132]. This architecture has emerged as a unifying framework across the three cornerstone areas of CV: perception, generation, and reasoning. A synergistic relationship has become evident, where advancements in one area often precipitate progress in the others, as illustrated in Figure 1. For instance, generative pretraining methods like Masked Image Modeling (MIM) [50, 6] have significantly boosted the performance of downstream perception tasks. Conversely, perception-centric features from models like DINOv2 [108] have been leveraged to guide and accelerate generative diffusion model training [145]. Such interplay is also demonstrated by the modern Vision-Language Models (VLMs) [59, 3, 21], ImageGen-CoT [81] and LLaDa [106]. However, a crucial link remains underexplored: the top-down influence of high-level reasoning on the foundational learning of low-level perceptual features. This raises a compelling question: can abstract reasoning, which often demands highly discriminative visual features, be harnessed to guide a perception model toward learning more precise and efficient representations?
Figure 2: Comparison of pretraining paradigms for ViT foundation models. ViTP employs an instruction-following objective to directly instil domain-specific perception capabilities into the vision backbone.
Answering this question is particularly critical for pretraining powerful foundation models in specialized domains like remote sensing and medical imaging, where prevailing paradigms face formidable challenges. Prevailing pretraining paradigms include supervised classification, MIM [50, 6, 112], image-image contrastive learning [49, 108, 41], and image-text contrastive learning [110, 88, 171]. Each approach, however, presents notable drawbacks. Supervised pretraining risks overfitting to narrow label distributions, thereby limiting generalization. MIM, by focusing on pixel-level reconstruction, may inadvertently neglect the fine-grained details of small but critical objects, a common scenario in remote sensing and medical scans. Image-image contrastive methods, while effective, are often notoriously difficult to optimize and demand substantial computational resources. These paradigms are all fundamentally “bottom-up”: they operate on the premise that robust low-level perception is a prerequisite for high-level reasoning [93, 84]. While this hierarchy is well-established, the reciprocal pathway (i.e. whether high-level reasoning can directly enhance low-level feature learning) remains largely unexplored.
Image-to-text contrastive learning, exemplified by CLIP [110], represents an initial step toward reasoning-guided pretraining. However, these methods primarily aim for general-purpose image-text alignment rather than optimizing a vision backbone for specialized, fine-grained downstream tasks. Consequently, the global, image-level features learned by models like CLIP’s ViT often prove suboptimal for dense prediction tasks such as semantic segmentation or object detection, which demand precise, pixel-level understanding [79]. This limitation underscores that merely aligning global image and text representations is an insufficient strategy for forging a powerful, domain-specific perception backbone.
Figure 3: ViTP sets new SOTA performance across a diverse range of downstream tasks in medical imaging and remote sensing.
To bridge this gap, we introduce Visual insTruction Pretraining (ViTP), a novel, top-down pretraining paradigm that directly integrates high-level, instruction-based reasoning into the perceptual feature learning process of a ViT backbone. As illustrated in Figure 2, ViTP diverges from prior methods by embedding the ViT within a larger Vision-Language Model (VLM) and pretraining it via a visual instruction-following objective. We construct our training data with image-text pairs derived directly from the target downstream tasks. During training, image tokens from the ViT and text tokens from a user query are fed jointly into a Large Language Model (LLM) [89]. The LLM’s generated response serves as the supervisory signal, compelling the ViT backbone to learn the complex data distributions of the downstream domain in an end-to-end fashion. This process endows the ViT with highly relevant, domain-specific perceptual capabilities. Furthermore, we introduce Visual Robustness Learning (VRL) by randomly dropping a large fraction of the ViT’s output image tokens before they are passed to the LLM. This constraint implicitly forces the ViT’s attention mechanism to encode more comprehensive and robust information within each of the remaining tokens, thereby enhancing the robustness and semantic richness of the learned visual features.
Extensive experiments on 16 challenging remote sensing and medical imaging benchmarks validate the effectiveness and efficiency of our approach. As shown in Figure 3, ViTP achieves new state-of-the-art results across several tasks. Notably, the ViTP pretraining process is computationally efficient, requiring only one day on 8 A40 GPUs. This work not only presents a novel pretraining paradigm but also offers a promising solution for creating powerful, domain-adapted foundation models.
Our contributions are summarized as follows:
  • We introduce ViTP, a top-down pretraining paradigm that leverages instruction following objectives to imbue a ViT backbone with high-level semantic perception tailored for specific downstream domains.
  • We propose Visual Robustness Learning, a regularization method that encourages the ViT to learn more comprehensive and robust feature representations by operating on a sparse set of visual tokens.
  • Extensive experiments demonstrate the pretraining efficiency and state-of-the-art performance of ViTP on downstream tasks in remote sensing and medical imaging.
2 Related Work
2.1 Perception Foundation Model
2.1.1 General Domains
Early approaches predominantly relied on supervised pretraining over large-scale labeled datasets such as ImageNet [47, 34]. While effective in learning semantically meaningful representations, these methods risk overfitting to the label space of the pretraining task, thereby limiting generalization to diverse downstream applications. Consequently, the field has increasingly shifted toward unsupervised learning paradigms that leverage unlabeled data. Among these, contrastive learning and masked image modeling (MIM) emerged as two dominant strategies. Contrastive learning methods, including MoCo [49], BYOL [39], and DINOv2 [108], learn discriminative representations by minimizing the distance between positive pairs (e.g., augmented views of the same image) while maximizing the distance between negative pairs in the embedding space. These methods emphasize semantic invariance and shown strong generalization capabilities across tasks. On the other hand, MIM-based approaches such as MAE [50], SimMIM [150] and UM-MAE [71] adopt a generative pretext task: a large portion of the input image is masked, and the model is trained to reconstruct the missing content. This encourages the model to learn holistic, contextually rich visual features by capturing spatial dependencies and fine-grained details.
While both paradigms achieved impressive results on natural images, their direct transfer to specialized domains such as remote sensing or medical imagery often results in suboptimal performance [27, 88, 42]. This underscores the necessity for domain-specific adaptations in pretraining strategies.
2.1.2 Remote sensing
In the remote sensing domain, pretraining strategies are tailored to address unique challenges such as large scale variations, multi-spectral imagery, and domain-specific object categories [41, 112, 68, 119]. Supervised methods like SAMRS [136] and MSFA [77] achieve strong performance by leveraging annotated remote sensing datasets. To exploit the abundance of unlabeled data, MIM-based approaches such as RingMo [122], SatMAE [27], and Scale-MAE [112] are proposed, specifically designed to model the dense small objects, multi-modal or multi-scale characteristics of satellite imagery. Meanwhile, contrastive learning has been explored through both image-text and image-image formulations. Models like GeoRSCLIP [171] and RemoteCLIP [88] leverage large-scale image-text pairs to enable zero-shot classification capabilities. Image-based contrastive methods such as CACo [101] and Skysense [41] focus on learning discriminative features from unlabeled remote sensing imagery. Hybrid approaches like CMID [104] and GFM [102] further integrate the strengths of both MIM and contrastive learning to learn more comprehensive and robust representations.
2.1.3 Medical Images
Pretraining vision foundation models is particularly crucial in medical imaging, where labeled data is scarce and expensive to obtain. A prominent line of research focuses on adapting the general-purpose Segment Anything Model (SAM) [60] to the medical domain. Adaptations range from finetuning SAM on large-scale medical datasets, as in MedSAM [100], to pretraining SAM-like architectures from scratch on medical images, as demonstrated by SAM-Med2D [24]. Other frameworks, such as IMIS-SAM [26], aim to enhance the interactive capabilities of SAM for clinical workflows. Concurrently, specialized MIM pretraining methods like MedMAE [42] and S3D [134] are developed to better capture the unique characteristics of medical data.
A common limitation across these domain-specific foundation models is their heavy reliance on tailored architectural or training designs, which restricts their transferability to other domains. In contrast, our proposed ViTP framework offers a more generalizable pretraining strategy that can be seamlessly adapted to various downstream domains by simply curating the corresponding visual instruction datasets.
2.2 Continual Pretraining
Continual pretraining was first popularized in NLP, where Gururangan et al. [43] demonstrated that continuing pretraining on in-domain data significantly improves model performance. This concept is also successfully adapted to computer vision. For instance, CSPT [168] first pretrains on ImageNet and then continues pretraining on a target remote sensing dataset using an MIM objective. TOV [126] adopts a curriculum strategy, freezing early layers pretrained on natural images while finetuning deeper layers on specialized data. To reduce computational overhead, RemoteCLIP [88] initializes from pretrained CLIP [110] weights and continues pretraining on domain-specific data. Similarly, MSFA [77] adopts a multi-stage strategy: pretraining on ImageNet, followed by training on optical remote sensing detection datasets, and finally finetuning on SAR detection data.
Inspired by these successes, we adopt a continual pretraining approach, initializing our VLM from the well-established InternVL [19] model. However, our method differs from prior work in two key aspects: (1) to mitigate catastrophic forgetting of general visual knowledge, we incorporate a small proportion of general-domain data into the specialized training corpus; (2) we introduce Visual Robustness Learning as an additional regularizer to enhance the model’s robustness during continual pretraining.
2.3 Visual Instruction Tuning
The rise of powerful VLMs such as Gemini [127], InternVL [20], and Qwen-VL [3] are largely driven by visual instruction tuning [89]. This paradigm aims to align visual representations with a pretrained Large Language Model (LLM) by projecting image features into the LLM’s embedding space. The model is then finetuned on a corpus of image-text instruction pairs, enabling the LLM to interpret visual inputs and respond to textual instructions. In this setup, the vision encoder typically remains frozen, serving solely as a feature extractor, while all learning occurs in the LLM and the projection module.
While this paradigm effectively leverages perception to enhance reasoning, the reverse direction, i.e. using reasoning to improve perception, remains largely underexplored. Our proposed ViTP addresses this gap by inverting the traditional visual instruction tuning pipeline. Instead of using images to tune an LLM, we utilize an LLM to guide the pretraining of the vision encoder.
3 Method
Figure 4: A conceptual illustration of the ViTP framework. A ViT backbone is embedded within a large VLM and then pretrained with domain-specific instruction following objective and Visual Robustness Learning (VRL). This process instils high-level semantic understanding into the ViT. The resulting weights are then used to initialize models for various downstream perception tasks.
In this section, we present our visual instruction pretraining (ViTP) framework, a novel paradigm designed to pretrain a ViT backbone. As illustrated in Figure 2, ViTP fundamentally differs from traditional pretraining by leveraging the reasoning capabilities of modern VLMs. The entire process is driven by a “visual instruction following objective”, where the ViT learns to extract features that help an LLM answer questions about an image. To tailor the model for specific downstream applications, we outline a “data recipe” for curating a domain-specific pretraining dataset. During pretraining, we employ “Visual Robustness Learning (VRL)”, a regularization technique that drops image tokens to enhance the semantic robustness of the learned representations. The final pretrained ViT serves as a powerful foundation model, readily adaptable to a variety of downstream tasks.
3.1 Visual Instruction Following Objective
The central hypothesis of ViTP is that a ViT can learn more potent and relevant features if its training is guided by high-level reasoning. To achieve this, we frame the pretraining as a visual instruction-following task. The framework, depicted in Figure 4, processes a domain-specific image through a ViT encoder to produce a sequence of image tokens. These tokens are projected into the LLM’s embedding space and concatenated with the tokenized text of an instruction. A Large Language Model (LLM) then processes this combined sequence to generate a response. The entire model is trained end-to-end, allowing the supervisory signal from the LLM’s response to directly optimize the ViT’s feature extraction process.
Our approach employs a continual pretraining strategy [43], starting with a well-trained, general-purpose VLM. We then continue its training on our curated domain-specific datasets. This offers two primary advantages: 1) It leverages the VLM’s vast pre-existing knowledge of general visual and linguistic patterns, providing a robust initialization. 2) This strong starting point significantly accelerates convergence during domain-specific pretraining, enhancing computational efficiency. Let the raw dataset be $\mathcal{D}_{\text{raw}}=\{(I,Q,R)\}$, where $I$ is an image, $Q$ is a text query (instruction), and $R$ is the ground-truth text response. This is processed into the final training set $\mathcal{D}=\{(x_{i},x_{t},y^{\star})\}$, where $x_{i}$, $x_{t}$, and $y^{\star}$ represent the processed image tokens, text tokens, and target response tokens, respectively.
3.1.1 ViT Feature Extraction and Projection
Given an input image $I\in\mathbb{R}^{H\times W\times 3}$, the ViT backbone partitions it into a grid of non-overlapping patches. Each patch is linearly embedded, and these patch embeddings are prepended with a ‘[CLS]’ token and processed through a series of Transformer blocks [132]. This yields a sequence of output image tokens $x^{\prime}_{i}=\{t_{1},t_{2},\ldots,t_{N}\}$, where $N$ is the sequence length. To align these visual tokens with the LLM’s embedding space, a lightweight projection layer (e.g., a two-layer MLP) maps these visual tokens $x^{\prime}_{i}$ into the final image tokens $x_{i}$.
3.1.2 Instruction-following Token Concatenation
For each image, the corresponding text query $Q$ is converted into a sequence of text tokens $x_{t}$ using the LLM’s tokenizer. These tokens represent the task-specific instruction. The projected image tokens $x_{i}$ and text tokens $x_{t}$ are then concatenated to form a unified input sequence. Crucially, learnable positional encodings are added to the embeddings of the image and text tokens to provide the LLM with spatial and sequential context. The final input sequence for the LLM is formed as:
$S_{llm}=[\text{PE}(x_{i})\,;\,\text{PE}(x_{t})]$ (1)
where $\text{PE}(\cdot)$ denotes the addition of positional encodings to the token embeddings and ‘$[;]$’ signifies sequence concatenation.
3.1.3 LLM-based Supervision
The combined sequence $S_{llm}$ is processed by the LLM, which acts as a reasoning engine to interpret the visual information from $x_{i}$ in the context of the instruction $x_{t}$. The model then auto-regressively generates an output sequence $O$. The entire system is optimized by minimizing the discrepancy between the generated output $O$ and the ground-truth target response $y^{\star}$. For instance, if the instruction asks to identify an object, $y^{\star}$ could be a textual description including its location in a structured format. During pretraining, the gradients from the output loss propagate back through the entire model, including the projection layer and the ViT backbone. We allow the weights of the ViT, the projection layer, and the LLM to be trainable. The optimization follows a standard supervised finetuning (SFT) objective, which minimizes the negative log-likelihood of the target sequence:
$\mathcal{L}_{\mathrm{SFT}}(\theta)=\mathbb{E}\Bigl[-\log P_{\theta}\bigl(y^{\star}\,\big|\,S_{llm}\bigr)\Bigr],$ (2)
where $(x_{i},x_{t},y^{\star})\sim\mathcal{D}$ and $P_{\theta}$ is the probability distribution over the text sequences parameterized by the entire model $\theta$.
3.2 Visual Robustness Learning
To foster the learning of more robust and semantically rich features, we introduce Visual Robustness Learning (VRL), a simple yet effective regularization technique applied during pretraining. As shown in Figure 4, VRL randomly drops a significant fraction of the projected image tokens $x_{i}$ before they are concatenated with the text tokens $x_{t}$. This operation is performed after positional encodings are associated with the tokens, ensuring the LLM retains knowledge of the original spatial positions of the surviving tokens. The VRL objective is thus:
$\mathcal{L}_{\mathrm{VRL}}(\theta)=\mathbb{E}\Bigl[-\log P_{\theta}\bigl(y^{\star}\,\big|\,[\mathcal{C}_{r}(\text{PE}(x_{i}));\text{PE}(x_{t})]\bigr)\Bigr],$ (3)
where $\mathcal{C}_{r}$ is a random sampling operation that drops a proportion $r$ of the tokens from a sequence. Formally, for a sequence $\mathcal{S}$:
$\mathcal{C}_{r}(\mathcal{S})\sim\bigl\{X\subseteq\mathcal{S}\;\big|\;|X|=\lceil(1-r)\cdot|\mathcal{S}|\rceil\bigr\},$ (4)
where the subset $X$ is selected uniformly at random while preserving the original ordering, and $r\in[0,1)$ . This “torture” mechanism forces the ViT to encode more comprehensive information in each token, as the model must infer the full visual context from a partial input. It encourages the ViT’s attention mechanism to learn robust, distributed and less redundant representations. As a practical benefit, dropping a large portion of tokens (e.g., $r=0.75$) significantly reduces memory usage and accelerates computation, enhancing ViTP’s scalability.
3.3 Pretraining Dataset Recipe
With the rapid development of VLMs, numerous image-text paired instruction datasets tailored for downstream domains have been released. The efficacy of ViTP is heavily dependent on the quality and composition of the pretraining dataset. We establish four key principles for constructing our data mixture:
  1. 1.
    Scale and Diversity: The dataset must be large and diverse, containing a wide array of visual concepts, scenes, and objects representative of the target domain.
  2. 2.
    Modality Coverage: The data must encompass all imaging modalities expected in downstream tasks. For instance, if a downstream task uses remote sensing Synthetic Aperture Radar (SAR) imagery, the pretraining mix should include such modality data to ensure the model learns modality-specific features.
  3. 3.
    Task Capability Alignment: The instruction-following tasks in the pretraining data should foster abilities required downstream. For instance, for downstream object detection tasks, including visual grounding and fine-grained VQA during pretraining empowers the backbone with localization and spatial reasoning capabilities.
  4. 4.
    Preservation of Generality: A certain fraction of general-domain natural images (e.g., from public VLM datasets) should be included. Domain-specific data can be limited in diversity, adding general data mitigates overfitting and prevents the model from losing its foundational ability to understand broad visual patterns.
3.4 Downstream Finetuning
Once ViTP is complete, the pretrained ViT backbone is then extracted to serve as a powerful backbone network. For downstream applications, this pretrained ViT is integrated into a standard ViT-Adapter [18] and combined with task-specific heads for detection, segmentation, or other tasks. The entire model is then finetuned on the target downstream dataset. This transfer learning approach leverages the rich, instruction-aware representations learned during ViTP, enabling faster adaptation and superior performance on specialized tasks.
Table 1: Overview of downstream task datasets for remote sensing, including object detection (horizontal and oriented), segmentation (semantic and instance), and change detection.
Task Dataset Modal Box Format Images Classes
  Object Detection DIOR [63] RGB Hori. Box 23,463 20
DIOR-R [25] RGB Ori. Box 23,463 20
DOTA-v2.0 [146] RGB Ori. Box 11,268 18
SARDet-100K [77] SAR Hori. Box 116,598 6
SSDD [167] SAR Hori. Box 1,160 1
RSAR [170] SAR Ori. Box 95,842 6
Semantic and Instance Segmentation iSAID [143] RGB Mask 2,806 15
LoveDA [139] RGB Mask 5,987 7
UAVid [99] RGB Mask 5,510 8
SSDD [167] SAR Polygons 1,160 1
Change Detection SVCD [62] RGB Mask 16,000 2
WHU [57] RGB Mask 11,456 2
LEVIR-CD [11] RGB Mask 10,192 2
S2Looking [115] RGB Mask 5,000 2
Medical Segmentation AMOS2022 [58] CT Mask 19,310 14
BraTS2021 [30] MRI Mask 76,467 3
CovidQUEx [124] X-ray Mask 5,826 2
3.5 Implementation Details
Training a VLM from scratch conventionally requires a multi-stage pipeline: vision–language contrastive pretraining, projector alignment, and large-scale instruction tuning. To bypass the prohibitive cost of these initial phases, we instead bootstrap from a publicly available, high-capacity VLM. Specifically, we initialise our weights from InternVL-2.5 [19], which uses a custmized ViT-Large backbone [20] and a Qwen2 [2] language model. All pretraining are performed on 8× NVIDIA A40 (48 GB) GPUs with a global batch size of 128, and all downstream task finetunings are conducted on $8\times$ NVIDIA RTX3090 (24 GB) GPUs. For ViTP pretraining, we use the AdamW optimizer with a learning rate of 2e-5 and a cosine decay schedule over 8,000 training steps. To accelerate both pre-training and downstream finetuning, Flash-Attention [29] is integrated into every ViT self-attention layer, yielding extra throughput gain without altering convergence behaviour.
4 Experiments
This section offers a rigorous assessment of ViTP’s finetuning capabilities on various downstream tasks from both remote-sensing and medical imaging domains. Dataset statistics and task specifications are summarized in Table 1. This section presents our experimental setup, a thorough analysis of the results, and an ablation study. We demonstrate ViTP’s superior performance compared to existing state-of-the-art methods and analyze its efficiency and robustness. The best score is indicated in bold, while the second-best score is underlined.
4.1 Remote Sensing
Table 2: Composition of the pretraining dataset for remote sensing.
       Dataset Size Sample Rate Tasks
  Mini-InternVL [38] 1394k 0.03 Caption,VQA,OCR
RSVQA [94] 100k 0.1 VQA
FIT_RS [98] 100k 0.1 VQA
GeoChat [61] 64k 2 VG
VRSBench [72] 38k 5 VG
RSVG [123] 5.5k 10 VG
DIOR-RSVG [162] 27k 8 VG
ISPRS_SAR [56] 1.5k 1 CLS
SAR_Sentinel-1&2 [131] 16k 1 CLS
VHM [109] 223k 1 Caption,VQA,CLS
LevirCCcaptions [87] 50k 0.5 Caption
GAIA [160] 33k 1 Caption
Million-AID [95] 920k 0.05 Caption,CLS
For the ViTP pretraining phase, we leverage a diverse set of publicly available image-text paired instruction datasets, adhering to the principles outlined in Section 3.3. The pretraining datasets are listed in Table 2. In the dataset recipe, Million-AID [95], GAIA [160], LevirCCcaptions [87], VHM [109], RSVQA [94], and FIT_RS [98] are large-scale remote sensing visual instruction datasets, providing diverse remote sensing scenes and visual tasks. ISPRS_SAR [56] and SAR Sentinel-1&2 [131] datasets specifically address the unique challenges of Synthetic Aperture Radar (SAR) data. We construct question-answer pairs related to object identification, scene interpretation, and attribute recognition in SAR images using these existing SAR classification datasets. This ensures modality alignment and task relevance for SAR-specific downstream applications. GeoChat [61], DIOR-RSVG [162], RSVG [123], and VRSBench [72] datasets primarily focus on visual grounding tasks within the remote sensing domain, providing rich instructional guidance for spatial reasoning, object attribute recognition, and target localization. The mini-InternVL [38] dataset contributes to general vision-language understanding capabilities, ensuring that ViTP retains broad applicability and does not overfit to highly specialized remote sensing patterns. The comprehensive pretraining dataset recipe enables ViTP to learn robust, instruction-aware visual representations that are highly relevant to diverse remote sensing tasks.
Table 3: Object detection performance (mAP %) on optical remote sensing datasets.
Model DIOR DIOR-R Model DOTA-v2
  GASSL [1] 67.40 65.65 RetinaNet [85] 46.68
SatMAE [27] 70.89 62.30 F-RCNN [113] 47.31
RingMo [122] 75.90 - FCOS [128] 48.51
CACO [101] 66.91 64.10 ATSS [166] 49.57
SSL4EO [142] 64.82 61.23 SASM [51] 44.53
CMID [104] 75.11 66.37 S2ANet [45] 49.86
RVSA [137] 73.22 70.96 KLD [156] 47.26
SatLas [7] 74.10 67.59 O-RepPoints [67] 48.95
GFM [102] 72.84 67.67 RoT Trans. [32] 52.81
ScaleMAE [112] 73.81 70.20 O-RCNN [149] 53.28
MA3E [80] - 71.82 GGHF [53] 57.17
Sel-MAE [138] 78.70 71.75 DCFL [151] 57.66
SkySense [41] 78.73 74.27 BillionFM [10] 58.69
ViTP 79.80 75.08 ViTP 60.23
4.1.1 Object detection
Object detection in remote sensing [146, 149, 78, 120, 177] involves identifying and precisely localizing objects of interest (e.g., vehicles, ships, and bridges) within aerial or satellite imagery. This task is crucial for applications such as urban planning and disaster monitoring. Challenges include vast scale variations of objects, arbitrary orientations, dense object distributions, and complex backgrounds. To evaluate the applicability of our proposed model for remote sensing object detection tasks under various scenarios, we conducted experiments on the following datasets: DIOR [63], DIOR-R [25], DOTA-v2.0 [146], SARDet-100K [77], RSAR [170], and SSDD [167]. We use the standard COCO evaluation metrics (mean Average Precision, mAP) to evaluate the performance of the models. Following previous practice [41, 170, 77], we use the oriented R-CNN [149] as the default detector for oriented object detection tasks and Cascade R-CNN [8] for horizontal object detection tasks. For training and testing, we resized all optical datasets (DIOR, DIOR-R, and DOTA-v2) to a standard size of $\times1024$ pixels. For the DOTA-v2 dataset specifically, we adopted a single-scale approach, cropping images into $\times1024$ patches with a 200-pixel overlap to handle its original large size. The SAR datasets, SARDet-100K and RSAR, were resized to $\times800$ pixels.
Table 4: Object detection performance (mAP %) on SAR datasets.
Model SARDet-100K Model RSAR
  DETR [9] 31.8 Def. DETR [178] 46.62
Sparse RCNN [121] 38.1 RetinaNet [85] 57.67
Dab-DETR [91] 45.9 ARS-DETR [161] 61.14
FCOS [128] 46.5 R3Det [155] 63.94
Grid RCNN [97] 48.8 LLMRotate [65] 64.1
GFL [70] 49.8 ReDet [46] 64.71
Defor. DETR [178] 50.0 O-RCNN [149] 64.82
MSFA [77] 53.7 S2ANet [45] 66.47
DenoDet [28] 55.4 RoI-Trans. [32] 66.95
DenoDetv2 [105] 56.4 SatMAE [27] 67.99
SARATR-X [69] 57.3 RemoteCLIP [88] 69.18
ViTP 59.7 ViTP 72.31
Table 3 presents the object detection results on RGB remote sensing datasets. ViTP consistently achieves state-of-the-art performance across DIOR, DIOR-R, and DOTA-v2.0. For DOTA-v2.0, which is particularly challenging due to a large quantity of small objects and dense scenes, ViTP achieves a new state-of-the-art of 60.23 mAP. This significantly surpasses previous top performers like BillionFM (58.69), highlighting ViTP’s superior ability to handle complex spatial relationships and arbitrary object orientations, likely due to its instruction-following pretraining that fosters fine-grained localization capabilities. To be noticed, SkySense [41] pretraining requires significantly more computational resources, over $17\times$ the GPU hours of ViTP, ViTP still outperforms SkySense. These results underscore the effectiveness of ViTP’s instruction-following pretraining in learning robust and task-relevant visual representations for optical remote sensing imagery.
Table 4 presents the object detection results on Synthetic Aperture Radar (SAR) datasets. SAR imagery poses unique challenges due to speckle noise, different scattering mechanisms, and a lack of visual texture compared to optical images. On SARDet-100K, ViTP achieves 59.7 mAP, significantly outperforming the previous state-of-the-art SARATR-X (57.3) and other detectors. For RSAR, ViTP sets a new state-of-the-art with 72.31 mAP. These remarkable improvements on SAR datasets demonstrate ViTP’s strong generalization capabilities and its effectiveness in handling challenging modalities.
Table 5: Semantic segmentation performance (mIoU %) on optical datasets.
model iSAID LoveDA model UAVid
  SeCo [157] 57.20 43.63 CANet [154] 63.50
DenseCLIP [111] 59.23 49.58 MP-Former [164] 63.67
SatMAE [27] 62.97 - ABCNet [66] 63.80
CACo [101] 64.32 48.89 DecoupleNet [96] 65.80
RVSA [137] 64.49 52.44 CoaT [153] 65.80
RSSFormer [152] 65.55 52.43 UNetFormer [141] 67.80
ScaleMAE [112] 65.77 - MaskFormer [23] 68.54
GASSL [1] 65.95 48.76 LSKNet [76] 70.00
CMID [104] 66.21 - Segmenter [117] 70.20
TOV [126] 66.24 49.70 RSSFormer [152] 70.69
RingMo [122] 67.20 - DeepLabv3+ [17] 71.33
SatLas [7] 68.71 - SegFormer [148] 71.44
Sel-MAE [138] - 53.92 DenseCLIP [111] 71.54
LSKNet [75] - 54.00 PSPNet [172] 71.71
SkySense [41] 70.91 - OCRNet [159] 71.84
ViTP 71.14 54.28 ViTP 73.39
4.1.2 Semantic Segmentation
Semantic segmentation in remote sensing [143, 141, 167] involves classifying each pixel in an image into a predefined category, such as land cover types (e.g., forest, water, urban area) or specific objects. This pixel-level understanding is vital for environmental monitoring, urban planning, and resource management. We evaluated ViTP’s performance on segmentation tasks using the following widely-used remote sensing datasets: iSAID [143], LoveDA [139], UAVid [99], and SSDD [167]. Following previous practice [41, 122], we use the UperNet [147] as the default segmentor for semantic segmentation tasks (iSAID, LoveDA, and UAVid) and Mask-RCNN [48] for instance segmentation tasks (SSDD). We report the mean Intersection over Union (mIoU) as the evaluation metric for semantic segmentation and Average Precision (AP) for instance segmentation tasks.
Table 5 and Table 6 demonstrate ViTP’s strong performance in remote sensing semantic segmentation. It sets new state-of-the-art performance on iSAID, LoveDA, UAVid, and SSDD datasets. Especially for UAVid and SSDD datasets, ViTP surpasses previous state-of-the-art by a large margin. The consistent improvements across these semantic segmentation benchmarks confirm that ViTP’s instruction-following pretraining on complex fine-grained regional understanding questions enables the ViT backbone to capture rich semantic information, which is crucial for accurate pixel-level classification.
Table 6: Detailed object detection and instance segmentation performance on the SSDD Dataset (SAR modality).
Model $AP_{box}$ $AP_{box}^{75}$ $AP_{box}^{75}$ $AP_{mask}$ $AP_{mask}^{50}$ $AP_{mask}^{75}$
  BoxInst [130] 44.76 83.75 44.11 34.10 71.16 27.27
Mask2Former [22] 53.40 78.45 67.02 56.52 85.10 69.48
InstaBoost [35] 54.77 87.85 58.54 58.95 89.05 71.57
CondInst [129] 57.89 92.53 67.40 50.31 90.46 54.80
SAM-Seg [60] 62.41 94.32 75.38 59.46 92.79 72.17
CATNet [92] 64.66 96.46 79.81 64.11 96.35 77.87
HQ-ISNet [118] 65.58 95.48 80.76 64.75 95.26 81.70
RSP-Query [16] 66.50 95.80 81.81 64.57 95.97 81.67
SCNet [133] 67.25 95.75 83.38 62.66 94.75 76.53
ViTP 70.80 97.80 86.60 65.90 96.80 81.80
Table 7: Change detection performance (F1-Score %) on SVCD, WHU, LEVER-CD and S2Looking datasets.
Model SVCD WHU LEVIR S2Looking
  Scale-MAE [112] - - 86.60 50.20
SeCo [157] - - 88.40 66.00
CACo [101] - - 89.20 65.90
GASSL [1] - - 89.60 66.30
SatMAE [27] - - 90.00 65.00
SatMAE++ [107] - - 90.70 56.40
CGNet [44] - - 92.01 64.33
Changer [37] - - 92.06 67.08
DiFormer [82] - - 92.15 66.31
Changen2 [176] - - 92.20 69.10
SkySense [41] - - 92.58 -
CLNet [175] 92.10 - 90.00 -
SRCDNet [90] 92.94 87.40 - -
ESCNet [163] 93.54 - - -
DSAMNet [116] 93.69 - - -
GCD-DDPM [144] 94.93 92.54 90.96 -
CDContrast [140] 95.11 - - -
DDPM-CD [5] 95.62 92.65 90.91 -
DMNet [73] 95.93 - - -
SNUnet [36] 96.20 83.49 88.59 63.19
BIT [12] - 83.98 89.31 63.76
BiFA [36] - 94.37 90.69 -
SGSLN [173] 96.24 94.67 91.93 -
RSP [135] 96.81 - 90.93 -
SAAN [40] 97.03 - 91.41 -
SiamixFormer [103] 97.13 - 91.58 -
TransUNetCD [64] 97.17 93.59 91.11 -
RDPNet [13] 97.20 - 91.20 -
WNet [125] 97.56 91.25 90.67 -
ChangeMamba [14] - 92.55 90.16 -
RS-Mamba [174] - 92.79 89.77 -
ChangeFormer [4] - 93.04 91.11 63.39
CDMamba [165] - 93.76 90.75 67.08
LSKNet [76] - 92.06 92.27 67.52
RVSA [137] 97.78 94.07 92.52 -
ChangeCLIP [33] 97.89 94.82 92.01 -
P2V-CD [83] 98.42 92.38 91.94 -
ViTP 98.63 94.98 92.67 69.89
4.1.3 Change Detection
Change detection in remote sensing focuses on identifying and characterizing differences in the state of an object or phenomenon by observing it at different times. This bitemporal analysis is fundamental for applications such as urban sprawl monitoring, deforestation tracking, and urbanization analysis. The primary challenges in this task include handling variations in illumination and atmospheric conditions between image acquisitions, precise image co-registration, and distinguishing meaningful semantic changes from irrelevant ones. To assess ViTP’s capabilities in this domain, we conduct experiments on three widely-recognized public datasets: SVCD [62], LEVIR-CD [11] and WHU-CD [57]. Following the common practice, we employ a simple Siamese UperNet-based framework [147] as the default detector and report the F1-Score as our primary evaluation metric.
The comprehensive results, presented in Table 7, show that ViTP achieves state-of-the-art performance across all of SVCD, WHU, LEVIR-CD and S2Looking datasets. For the SVCD dataset, which contains more diverse change types, ViTP shows a clear performance advantage. On the LEVIR-CD dataset, which focuses on building changes, ViTP achieves a new state-of-the-art F1 of 92.67. This surpasses previous methods, demonstrating its proficiency in identifying fine-grained changes in complex suburban environments. Similarly, on the WHU-CD dataset, ViTP sets a new benchmark with an F1 of 94.98, effectively handling variations in building scale and appearance. These significant improvements underscore the rich semantic understanding power of ViTP for temporal analysis.
4.2 Medical Imaging
Table 8: Composition of the pretraining dataset for medical imagery.
       Dataset Size Sample Rate Tasks
  Mini-InternVL [38] 1394k 0.02 Caption,VQA,OCR
GMAI-MMBench [158] 5k 20 VQA
Open-i [31] 7k 5 Caption
Huatuo-OA [15] 647k 0.1 Caption
Huatuo-VQA [15] 647k 0.5 VQA
PMC-OA [86] 1647k 0.05 Caption
PMC-VQA [169] 227k 0.5 VQA
OmniMedical [52] 89k 1 VQA
Quilt-1M [54] 723k 0.1 Caption
Quilt-Instruct [114] 147 1 VQA
Similar to the approach for remote sensing, we collected several open-source medical image-text paired instruction datasets. The pretraining datasets are listed in Table 8. Among them, Huatuo [15], PMC [86], and OmniMedical [52] are large-scale, general medical visual instruction datasets covering various medical imaging modalities, including CT, MRI, ultrasound, and others. These datasets reframe various downstream tasks into a question-answering format, augmented with millions of descriptive captions, enabling the model to build comprehensive knowledge across diverse medical imaging modalities. Moreover, since medical image segmentation is a key downstream task, we also include the GAMI-MMBench [158] dataset. It provides VQA based on ROIs and segmentation masks, which aids in aligning the model with dense prediction tasks. The Open-i [31] and Quilt [54] datasets are included to compensate for the relative scarcity of X-ray and pathology images in other general medical collections. Finally, as with the remote sensing domain, mini-InternVL [38] is incorporated to enhance general vision-language understanding.
4.2.1 Semantic Segmentation
Semantic segmentation in medical image analysis involves classifying each pixel in an image into a predefined category, such as an organ (e.g., lung, kidney, heart) or a specific lesion. This pixel-level understanding is vital for clinical diagnosis, disease analysis, and patient prognosis. We evaluated ViTP’s performance on segmentation tasks using three challenging medical datasets: AMOS2022 [58], BraTS2021 [30], and ConvidQUEx [124]. These datasets encompass three primary clinical modalities: CT (AMOS2022), MRI (BraTS2021), and chest X-ray (ConvidQUEx), and covering both organ and lesion segmentation scenarios. We use UperNet [147] as the default segmentation head. We report the mean Dice Score (mDice) as the evaluation metric for all segmentation tasks.
Table 9 demonstrates the strong performance of ViTP in medical semantic segmentation. ViTP achieves new state-of-the-art results on the AMOS, BraTS, and ConvidQUEx datasets. Specifically, similar to ViTP, SAM-based approaches (MedSAM [100], SAMMed2D [24] and IMIS-SAM [26]) are also finetuned on the downstream datasets. However, ViTP achieves substantial performance gains, which can be attributed to its more effective pretraining strategy. When compared to the classic medical segmentation model (nnU-Net [55]), ViTP not only exhibits superior generalization but also surpasses its performance, despite nnU-Net being a highly specialized model. Finally, a comparison with MedMAE [42], a recent 2D medical pretraining method, highlights the effectiveness of our proposed pretraining paradigm.
Table 9: Semantic segmentation performance (mDice %) on Medical Imagery Datasets.
model AMOS2022 BraTS2021 CovidQUEx
  5-points Prompt
MedSAM [100] 81.36 69.44 78.18
SAMMed2D [24] 87.81 70.70 83.39
IMIS-SAM [26] 87.42 64.77 86.28
Box Prompt
MedSAM [100] 86.74 71.23 78.87
SAMMed2D [24] 88.67 71.26 77.81
IMIS-SAM [26] 88.71 70.59 82.91
nnUNet [55] 87.28 71.03 90.41
MedMAE [42] 90.11 69.59 90.18
ViTP 90.60 72.11 91.55
5 Analysis and Ablation Study
To comprehensively understand the impact of various components and hyperparameters on ViTP’s performance, we conduct a series of rigorous ablation studies and analytical experiments. This section details the ablation studies, including: (1) validating the principles of our data recipe, (2) assessing the impact of pre-training steps, (3) evaluating the effect of the VRL image token dropping ratio, and (4) analyzing the influence of the ViTP language model’s size.
For these studies, we utilize the ViTP model pre-trained on remote sensing data and evaluate its performance on the RSAR dataset, a challenging benchmark for oriented object detection in SAR imagery. We selected this specific downstream task for several reasons. Firstly, the SAR modality presents unique challenges not found in optical imagery, making it an excellent testbed for a model’s ability to generalize to specialized visual domains. Secondly, oriented object detection demands precise localization, which stringently tests the model’s capacity to capture fine-grained spatial information and recognize weak patterns amidst dense objects. Therefore, performance on the RSAR dataset serves as a fair and insightful proxy for the model’s overall representational power. For experimental efficiency, unless otherwise specified, models are finetuned on the RSAR validation set and evaluated on its test set.
Furthermore, we investigate the unique advantages of the ViTP paradigm over other pre-training methods during the finetuning stage, specifically focusing on its data efficiency and robustness to various data corruptions.
Table 10: Validation of the data recipe. Performance on the RSAR Benchmark confirms the necessity of our comprehensive data curation strategy.
Pretrain Paradigm RSAR mAP
  w/o Diversity 52.6
w/o SAR 52.5
w/o Grounding 53.0
w/o General data 52.3
full data 54.6
5.1 Data Recipe
Our ViTP pretraining dataset was carefully curated according to the principles stated in Section 3.3. Table 10 ablates the contribution of different data types, where “w/o Diversity” refers to the exclusion of diverse remote sensing visual instruction datasets (Million-AID, GAIA, Levir-CC, VHM, VRSBench, and RSVG) during pre-training. Removing any single component: diverse datasets, SAR-specific datasets, grounding datasets, or general-domain datasets, leads to a significant drop in downstream performance on the RSAR dataset. Excluding SAR data prevents the model from learning modality-specific features, while removing grounding data impairs its localization ability. Omitting general data leads to overfitting on specialized patterns, which leads to the most severe performance drop. This confirms that a diverse mixture of diverse, general, domain-specific, and task-specific instruction data is crucial for optimal performance.
5.2 Impact of Pretraining Steps
Figure 5: Effect of Pretraining Duration. RSAR mAP improves with more pretraining steps before saturating at approximately 8k steps.
The duration of pre-training is a significant factor influencing downstream task performance. As depicted in Figure 5, the model’s mAP on the RSAR dataset generally improves with an increasing number of pre-training steps. This trend indicates that more extensive instruction tuning enables the model to learn progressively better feature representations. However, the performance curve begins to plateau around 8k steps, with no further improvements observed thereafter. Consequently, we set the number of ViTP pre-training steps to 8k by default for our main model, balancing performance with computational cost. This pre-training phase takes approximately 23 hours to complete 8k steps training on $8\times$ A40 GPUs, demonstrating the efficiency of our proposed method.
5.3 Impact of VRL Token drop Ratio
The Visual Robustness Learning (VRL), introduced in Section 3.2, involves randomly dropping a portion of image tokens during pre-training. This study investigates the optimal dropping ratio for this process. An appropriate ratio compels the ViT to learn more robust and semantically rich representations by inferring missing visual context. Conversely, an excessively high ratio risks obscuring essential image information, thereby posing an ill-posed training target and hindering effective learning.
As shown in Figure 7, the model’s performance on the RSAR dataset varies significantly with different dropping ratios. The performance peaks at a ratio of 75%, which improves the mAP from 52.8 (without VRL) to 54.6. This suggests that a 75% masking ratio effectively regularizes the model without excessively impeding information flow. Ratios from 0% to 50% appear to provide insufficient regularization. Surprisingly, even with a very high masking rate of 90%, the model still benefits from this strong regularization, achieving a notable 54.3 mAP.
Figure 6: Impact of the VRL Drop Ratio. ViTP performance on RSAR reaches a peak with a 75% token drop ratio.
Figure 7: Impact of LLM size. A larger LLM hinders the performance of ViTP on the RSAR Benchmark.
5.4 Impact of Language Model Size
This study investigates the influence of the Large Language Model’s (LLM) capacity on ViTP’s overall performance. As shown in Figure 7, as the LLM size increases, the downstream performance on the RSAR dataset steadily degrades. This phenomenon is analogous to the findings in Masked Autoencoders [50] pretrain paradigm, which highlight the importance of a lightweight decoder. It indicates that for the challenging SAR modality, a more powerful LLM might compensate for suboptimal visual features extracted by the Vision Transformer (ViT). This could reduce the optimization pressure on the ViT, preventing it from learning the most effective modality-specific representations.
Table 11: Pretraining efficiency and performance. Time is estimated for pretraining on $8\times$ A40 GPUs. ViTP achieves SOTA or competitive performance with a fraction of the computational cost.
Methods Hours DIOR-R iSAID
  MIM RVSA [137] 250 70.96 64.49
Scale-MAE [112] 60 70.20 65.77
CL RemoteCLIP [88] 100 70.20 62.53
SkySense [41] 400 74.27 70.91
ViTP 23 75.08 70.54
5.5 Training Efficiency
A key advantage of the ViTP paradigm is its exceptional pretraining efficiency, which lowers the barrier for developing specialized foundation models. As detailed in Table 11, ViTP pretraining completes in one day on 8 A40 GPUs. This is dramatically faster than competing paradigms. For instance, ViTP is 2.6x faster than the efficient MIM-based method (Scale-MAE) and over 17x faster than the high-performing but computationally intensive contrastive learning method, SkySense.
Table 12: Data efficiency on the RSAR Benchmark. ViTP demonstrates significantly stronger performance in low-data regimes compared to MIM and contrastive methods.
Model 100% 20% 10% 5% 2%
  RemoteCLIP [88] 69.18
63.14
↓6.04
57.10
↓12.08
47.25
↓21.93
34.78
↓34.40
SatMAE [27] 67.99
61.36
↓6.63
55.24
↓12.75
50.76
↓17.23
37.90
↓30.09
ViTP 72.31
67.07
$_{\downarrow\textbf{5.24}}$
61.68
$_{\downarrow\textbf{10.63}}$
56.42
$_{\downarrow\textbf{15.89}}$
46.98
$_{\downarrow\textbf{25.33}}$
Table 13: Model robustness to common image corruptions. ViTP demonstrates superior resilience compared to MIM and CL baselines on the REOBench benchmark (DIOR-R, mAP). It maintains the highest average performance across all corruptions and suffers the smallest performance drop ($\Delta_{\text{TP}}$) from the clean baseline. The inclusion of Visual Robustness Learning (VRL) further enhances this robustness.
   Model Clean Bright Cloud Comp. Data Gauss Gauss Haze Motion Rotate Salt Scale Trans Avg $\Delta_{\text{TP}}\downarrow$
Contrast Artifact Gaps Blur Noise Blur Pepper
  MIM SatMAE [27] 62.30 56.84 57.86 55.80 58.36 55.38 58.44 59.34 56.92 56.60 53.76 51.58 60.90 56.82 5.49
ScaleMAE [112] 70.20 64.80 65.98 62.50 64.46 62.58 63.82 66.10 63.08 63.44 60.50 53.08 68.26 63.22 6.98
RVSA [137] 70.96 60.59 65.02 61.58 64.60 62.35 62.87 63.98 62.88 64.04 56.61 55.97 69.69 62.51 8.45
SatMAE++ [107] 65.20 59.44 61.02 60.30 59.88 59.66 61.06 61.72 59.56 59.14 58.64 48.48 64.70 59.47 5.73
CL RemoteCLIP [88] 70.20 66.52 66.62 63.84 65.40 63.62 63.68 66.76 62.66 63.52 59.16 57.42 68.64 63.99 6.21
GeoRSCLIP [171] 69.80 66.12 65.34 65.34 64.96 63.62 62.90 66.04 62.02 62.68 56.04 57.40 68.10 63.38 6.42
ViTP w/o VRL 71.91 69.11 67.25 66.20 66.87 67.67 67.47 67.12 65.94 65.54 66.87 65.23 70.34 67.13 4.78
ViTP 73.37 70.56 69.54 67.74 70.58 69.05 68.97 70.85 67.86 67.10 67.23 66.64 71.87 69.00 4.37
5.6 Data Efficiency
We investigate ViTP’s performance in low-data regimes by finetuning it on subsets of the RSAR training dataset, ranging from 2% to 20%. To fully assess the modeling potential of each method, we extend the finetuning schedule to 36 epochs for this specific study. As shown in Table 12, ViTP consistently outperforms SatMAE[27] (an MIM method) and RemoteCLIP [88] (a contrastive method) across all data fractions. ViTP’s advantage becomes more pronounced as the amount of training data decreases. For instance, with only 2% of the training data, ViTP achieves 46.98 mAP, substantially outperforming SatMAE (37.90 mAP) and RemoteCLIP (34.78 mAP). Notably, ViTP finetuned on just 20% of the data (67.07 mAP) achieves performance comparable to or exceeding many state-of-the-art methods trained on the full dataset (as shown in Table 4). This superior data efficiency suggests that the rich semantic priors acquired during instruction tuning enable the model to generalize more effectively from a limited number of examples.
5.7 Model Robustness
While most research datasets contain curated clean imagery, real-world remote sensing images are often degraded by atmospheric conditions (e.g., clouds, haze), sensor noise, or processing artifacts. To evaluate ViTP’s resilience to such degradations, we assess its robustness against 12 common image corruptions from the REOBench [74] benchmark on the DIOR-R dataset. We retrained our ViTP model on the DIOR-R dataset using an image size of $\times800$ pixels and without any test-time augmentation, matching the settings of the REOBench baseline models. As shown in Table 13, the standard ViTP model is significantly more robust than all Masked Image Modeling (MIM) and Contrastive Learning (CL) models. It achieves the highest average mAP of 69.00 across all corruptions and the lowest performance drop ($\Delta_{\text{TP}}$) of 4.37 from its clean-image performance. We also ablate the effect of our proposed Visual Robustness Learning (VRL). The model trained without VRL already demonstrates superior robustness over the baselines. The inclusion of VRL further boosts the average performance from 67.13 to 69.00 mAP and reduces the performance degradation ($\Delta_{\text{TP}}$) from 4.78 to 4.37. This confirms that our instruction tuning paradigm, enhanced with VRL, is a highly effective technique for learning robust and comprehensive feature representations.
6 Limitation and Future Work
Despite the strong performance and efficiency of ViTP, we acknowledge several limitations that open promising avenues for future research. A primary limitation is that ViTP’s success is intrinsically linked to the quality and diversity of the instruction-following dataset. As our experiments suggest, a carefully curated data recipe is important for achieving optimal performance. The current process of collecting, filtering, and balancing datasets for specialized domains like remote sensing or medical imaging requires considerable manual effort and deep domain expertise. Future work could focus on developing more automated or semi-automated strategies for generating high-quality, domain-specific instruction data. Leveraging large language models to synthesize diverse question-answer or description-grounding pairs from unlabeled images could be a promising direction to enhance the scalability and accessibility of our pretraining paradigm. Moreover, our current work focuses on pretraining 2D image-based Vision Transformers. Extending the ViTP framework to other data modalities, such as video and 3D point clouds, represents another exciting frontier. Adapting the instruction-following objective to these domains could unlock new capabilities in temporal reasoning and 3D spatial understanding.
7 Conclusion
In this paper, we addressed the underexplored, top-down pathway from high-level reasoning to low-level perception in vision foundation models. We introduced Visual insTruction Pretraining (ViTP), a novel paradigm that leverages the high-level semantic reasoning of a Vision-Language Model to directly guide the feature learning of a ViT backbone. Complemented by our Visual Robustness Learning (VRL), ViTP compels the vision model to learn rich, robust, and task-relevant representations from downstream-specific instructional data. Our extensive experiments on 16 challenging remote sensing and medical imaging benchmarks validated the effectiveness of the proposed ViTP method. ViTP not only established new state-of-the-art performance across a diverse range of downstream tasks but also proved to be remarkably computationally efficient. We believe this work opens a promising new avenue for a deeper integration of high-level reasoning into the core of visual feature learning.

### v3

@@
 Abstract
-Modern computer vision is converging on a closed loop in which perception, reasoning and generation mutually reinforce each other.
-However, this loop remains incomplete: the top-down influence of high-level reasoning on the foundational learning of low-level perceptual features is not yet underexplored.
-This paper addresses this gap by proposing a new paradigm for pretraining foundation models in downstream domains.
-We introduce Visual insTruction Pretraining (ViTP), a novel approach that directly leverages reasoning to enhance perception.
-ViTP embeds a Vision Transformer (ViT) backbone within a Vision-Language Model and pretrains it end-to-end using a rich corpus of visual instruction data curated from target downstream domains.
-ViTP is powered by our proposed Visual Robustness Learning (VRL), which compels the ViT to learn robust and domain-relevant features from a sparse set of visual tokens.
+Human visual perception is the product of a synergistic interplay between bottom-up perception processing and top-down understanding-driven modulation.
+While modern computer vision has excelled at creating powerful models that mirror the bottom-up, hierarchical nature of perception, the reciprocal top-down influence of high-level understanding on the foundational learning of low-level perception features remains critically underexplored.
+This paper addresses this gap by proposing a new paradigm for pretraining foundation models, inspired by the top-down mechanisms of human vision.
+We introduce Visual insTruction Pretraining (ViTP), a novel approach that directly leverages understanding to enhance perception.
+ViTP embeds a Vision Transformer backbone within a Vision-Language Model and pretrains it end-to-end using a rich corpus of visual instruction data curated from target downstream domains.
+ViTP is powered by our proposed Visual Robustness Learning (VRL), which compels the Vision Transformer to learn robust and domain-relevant features from a sparse set of visual tokens.
 Extensive experiments on 16 challenging remote sensing and medical imaging benchmarks demonstrate that ViTP establishes new state-of-the-art performance across a diverse range of downstream tasks.
-The code is available at https://github.com/zcablii/ViTP.
+The code is available at https://github.com/zcablii/ViTP
 1 Introduction
-Figure 1: The synergistic relationship between perception, generation, and reasoning in modern CV.
-Our proposed ViTP forges a novel link from high-level reasoning to low-level perception, a previously underexplored connection.
-Recent years have witnessed a leap forward in computer vision (CV), largely catalyzed by the Transformer architecture [132].
-This architecture has emerged as a unifying framework across the three cornerstone areas of CV: perception, generation, and reasoning.
-A synergistic relationship has become evident, where advancements in one area often precipitate progress in the others, as illustrated in Figure 1.
-For instance, generative pretraining methods like Masked Image Modeling (MIM) [50, 6] have significantly boosted the performance of downstream perception tasks.
-Conversely, perception-centric features from models like DINOv2 [108] have been leveraged to guide and accelerate generative diffusion model training [145].
-Such interplay is also demonstrated by the modern Vision-Language Models (VLMs) [59, 3, 21], ImageGen-CoT [81] and LLaDa [106].
-However, a crucial link remains underexplored: the top-down influence of high-level reasoning on the foundational learning of low-level perceptual features.
-This raises a compelling question: can abstract reasoning, which often demands highly discriminative visual features, be harnessed to guide a perception model toward learning more precise and efficient representations?
-Figure 2: Comparison of pretraining paradigms for ViT foundation models.
+Decades of research in cognitive neuroscience have established that human visual perception is not a one-way street.
+It is a dynamic process arising from the constant interaction of two fundamental pathways: a bottom-up, data-driven pathway and a top-down, knowledge-driven one.
+The bottom-up process, famously elucidated by Hubel and Wiesel [62], involves a hierarchical construction of understanding: simple features like edges and orientations are extracted by early visual cortices and progressively assembled into complex object representations in higher cortical areas [117, 9].
+This paradigm has been the primary inspiration for modern computer vision.
+For instance, prevailing Vision Transformer pretraining paradigms include supervised classification, MIM [55, 7, 132], contrastive learning [56, 127, 49], and image-text contrastive learning [129, 102, 192].
+Recent multimodal large models [3, 25] usually utilize a vision tower to provide vision elements for high-level understanding.
+These paradigms are all fundamentally “bottom-up”: they operate on the premise that robust low-level perception is a prerequisite for high-level understanding [108, 97].
+Figure 1: (a) The synergistic relationship between perception and understanding in modern CV.
+Our proposed ViTP forges a previously underexplored link from high-level understanding to low-level perception.
+(b) Self-attention activation maps for the query patch (marked with a red cross).
+ViTP identifies fine-grained object parts that are high-level semantically related.
+However, this bottom-up hierarchy only tells half the story.
+A wealth of research in cognitive neuroscience has demonstrated the profound top-down influence of high-level understanding, attention, and prior knowledge on foundational perception.
+Cognitive processes, guided by expectation and goals, actively modulate and refine how low-level visual information is processed, even at the earliest stages in the visual cortex [70, 130, 123].
+This feedback mechanism allows the brain to focus on relevant details, fill in missing information, and interpret ambiguous stimuli with remarkable efficiency.
+While this reciprocal pathway—where understanding enhances perception—is integral to human vision, it remains critically underexplored in the computer vision models.
+This raises a compelling question: can abstract, high-level understanding be harnessed to directly guide a perception model toward learning more precise and domain-relevant representations?
+Figure 2: Comparison of pretraining paradigms for Vision Transformer (ViT) foundation models.
 ViTP employs an instruction-following objective to directly instil domain-specific perception capabilities into the vision backbone.
 Answering this question is particularly critical for pretraining powerful foundation models in specialized domains like remote sensing and medical imaging, where prevailing paradigms face formidable challenges.
-Prevailing pretraining paradigms include supervised classification, MIM [50, 6, 112], image-image contrastive learning [49, 108, 41], and image-text contrastive learning [110, 88, 171].
-Each approach, however, presents notable drawbacks.
-Supervised pretraining risks overfitting to narrow label distributions, thereby limiting generalization.
-MIM, by focusing on pixel-level reconstruction, may inadvertently neglect the fine-grained details of small but critical objects, a common scenario in remote sensing and medical scans.
-Image-image contrastive methods, while effective, are often notoriously difficult to optimize and demand substantial computational resources.
-These paradigms are all fundamentally “bottom-up”: they operate on the premise that robust low-level perception is a prerequisite for high-level reasoning [93, 84].
-While this hierarchy is well-established, the reciprocal pathway (i.e.
-whether high-level reasoning can directly enhance low-level feature learning) remains largely unexplored.
-Image-to-text contrastive learning, exemplified by CLIP [110], represents an initial step toward reasoning-guided pretraining.
+The supervised [152] pretraining risks overfitting to narrow label distributions; the MIM [55] may inadvertently neglect the fine-grained details of small but critical objects; image-image contrastive methods, while effective, are often notoriously difficult to optimize and demand substantial computational resources.
+Prevailing pretraining paradigms of image-to-text contrastive learning [129] or multimodal autoregressive Vision Transformer pretraining [44] represent steps toward reasoning-guided pretraining.
 However, these methods primarily aim for general-purpose image-text alignment rather than optimizing a vision backbone for specialized, fine-grained downstream tasks.
-Consequently, the global, image-level features learned by models like CLIP’s ViT often prove suboptimal for dense prediction tasks such as semantic segmentation or object detection, which demand precise, pixel-level understanding [79].
+Consequently, the global, image-level features learned by models like CLIP [129] or AIM [44] often prove suboptimal for dense prediction tasks such as semantic segmentation or object detection [89].
 This limitation underscores that merely aligning global image and text representations is an insufficient strategy for forging a powerful, domain-specific perception backbone.
-Figure 3: ViTP sets new SOTA performance across a diverse range of downstream tasks in medical imaging and remote sensing.
-To bridge this gap, we introduce Visual insTruction Pretraining (ViTP), a novel, top-down pretraining paradigm that directly integrates high-level, instruction-based reasoning into the perceptual feature learning process of a ViT backbone.
-As illustrated in Figure 2, ViTP diverges from prior methods by embedding the ViT within a larger Vision-Language Model (VLM) and pretraining it via a visual instruction-following objective.
+To bridge this gap, we introduce Visual insTruction Pretraining (ViTP), a novel, top-down pretraining paradigm that directly integrates high-level, instruction-based understanding into the perceptual feature learning process of a Vision Transformer backbone.
+As illustrated in Figure 2, ViTP diverges from prior methods by embedding the Vision Transformer within a larger Vision-Language Model (VLM) and pretraining it via a visual instruction-following objective.
 We construct our training data with image-text pairs derived directly from the target downstream tasks.
-During training, image tokens from the ViT and text tokens from a user query are fed jointly into a Large Language Model (LLM) [89].
-The LLM’s generated response serves as the supervisory signal, compelling the ViT backbone to learn the complex data distributions of the downstream domain in an end-to-end fashion.
-This process endows the ViT with highly relevant, domain-specific perceptual capabilities.
-Furthermore, we introduce Visual Robustness Learning (VRL) by randomly dropping a large fraction of the ViT’s output image tokens before they are passed to the LLM.
-This constraint implicitly forces the ViT’s attention mechanism to encode more comprehensive and robust information within each of the remaining tokens, thereby enhancing the robustness and semantic richness of the learned visual features.
+During training, image tokens from the Vision Transformer and text tokens from a user query are fed jointly into a Large Language Model (LLM) [103].
+The LLM’s generated response serves as the supervisory signal, compelling the Vision Transformer backbone to learn the complex data distributions of the downstream domain in an end-to-end fashion.
+This process endows the Vision Transformer with highly relevant, domain-specific perceptual capabilities.
+Furthermore, we introduce Visual Robustness Learning (VRL) by randomly dropping a large fraction of the Vision Transformer’s output image tokens before they are passed to the LLM.
+This constraint implicitly forces the Vision Transformer’s attention mechanism to encode more comprehensive and robust information within each of the remaining tokens, thereby enhancing the robustness and semantic richness of the learned visual features.
 Extensive experiments on 16 challenging remote sensing and medical imaging benchmarks validate the effectiveness and efficiency of our approach.
@@
 This work not only presents a novel pretraining paradigm but also offers a promising solution for creating powerful, domain-adapted foundation models.
+Figure 3: ViTP sets new SOTA performance across a diverse range of downstream tasks in medical imaging and remote sensing.
 Our contributions are summarized as follows:
-  • We introduce ViTP, a top-down pretraining paradigm that leverages instruction following objectives to imbue a ViT backbone with high-level semantic perception tailored for specific downstream domains.
-  • We propose Visual Robustness Learning, a regularization method that encourages the ViT to learn more comprehensive and robust feature representations by operating on a sparse set of visual tokens.
+  • We introduce ViTP, the first top-down pretraining paradigm that leverages understanding tasks to imbue a Vision Transformer backbone with high-level semantic perception , as illustrated in Figure 1.
+  • We propose Visual Robustness Learning, a regularization method that encourages the Vision Transformer to learn more comprehensive and robust feature representations by operating on a sparse set of visual tokens.
   • Extensive experiments demonstrate the pretraining efficiency and state-of-the-art performance of ViTP on downstream tasks in remote sensing and medical imaging.
@@
 2.1 Perception Foundation Model
-2.1.1 General Domains
 Early approaches predominantly relied on supervised pretraining over large-scale labeled datasets such as ImageNet [58, 40].
@@
 This underscores the necessity for domain-specific adaptations in pretraining strategies.
-2.1.2 Remote sensing
-In the remote sensing domain, pretraining strategies are tailored to address unique challenges such as large scale variations, multi-spectral imagery, and domain-specific object categories [41, 112, 68, 119].
-Supervised methods like SAMRS [136] and MSFA [77] achieve strong performance by leveraging annotated remote sensing datasets.
-To exploit the abundance of unlabeled data, MIM-based approaches such as RingMo [122], SatMAE [27], and Scale-MAE [112] are proposed, specifically designed to model the dense small objects, multi-modal or multi-scale characteristics of satellite imagery.
-Meanwhile, contrastive learning has been explored through both image-text and image-image formulations.
-Models like GeoRSCLIP [171] and RemoteCLIP [88] leverage large-scale image-text pairs to enable zero-shot classification capabilities.
-Image-based contrastive methods such as CACo [101] and Skysense [41] focus on learning discriminative features from unlabeled remote sensing imagery.
-Hybrid approaches like CMID [104] and GFM [102] further integrate the strengths of both MIM and contrastive learning to learn more comprehensive and robust representations.
-2.1.3 Medical Images
-Pretraining vision foundation models is particularly crucial in medical imaging, where labeled data is scarce and expensive to obtain.
-A prominent line of research focuses on adapting the general-purpose Segment Anything Model (SAM) [60] to the medical domain.
-Adaptations range from finetuning SAM on large-scale medical datasets, as in MedSAM [100], to pretraining SAM-like architectures from scratch on medical images, as demonstrated by SAM-Med2D [24].
-Other frameworks, such as IMIS-SAM [26], aim to enhance the interactive capabilities of SAM for clinical workflows.
-Concurrently, specialized MIM pretraining methods like MedMAE [42] and S3D [134] are developed to better capture the unique characteristics of medical data.
-A common limitation across these domain-specific foundation models is their heavy reliance on tailored architectural or training designs, which restricts their transferability to other domains.
-In contrast, our proposed ViTP framework offers a more generalizable pretraining strategy that can be seamlessly adapted to various downstream domains by simply curating the corresponding visual instruction datasets.
-2.2 Continual Pretraining
-Continual pretraining was first popularized in NLP, where Gururangan et al.
-[43] demonstrated that continuing pretraining on in-domain data significantly improves model performance.
-This concept is also successfully adapted to computer vision.
-For instance, CSPT [168] first pretrains on ImageNet and then continues pretraining on a target remote sensing dataset using an MIM objective.
-TOV [126] adopts a curriculum strategy, freezing early layers pretrained on natural images while finetuning deeper layers on specialized data.
-To reduce computational overhead, RemoteCLIP [88] initializes from pretrained CLIP [110] weights and continues pretraining on domain-specific data.
-Similarly, MSFA [77] adopts a multi-stage strategy: pretraining on ImageNet, followed by training on optical remote sensing detection datasets, and finally finetuning on SAR detection data.
-Inspired by these successes, we adopt a continual pretraining approach, initializing our VLM from the well-established InternVL [19] model.
-However, our method differs from prior work in two key aspects: (1) to mitigate catastrophic forgetting of general visual knowledge, we incorporate a small proportion of general-domain data into the specialized training corpus; (2) we introduce Visual Robustness Learning as an additional regularizer to enhance the model’s robustness during continual pretraining.
 2.2 Visual Instruction Tuning
 The rise of powerful VLMs such as Gemini [147], InternVL [25], and Qwen-VL [4] are largely driven by visual instruction tuning [103].
-This paradigm aims to align visual representations with a pretrained Large Language Model (LLM) by projecting image features into the LLM’s embedding space.
+This paradigm aims to align visual representations with a pretrained Large Language Model by projecting image features into the LLM’s embedding space.
 The model is then finetuned on a corpus of image-text instruction pairs, enabling the LLM to interpret visual inputs and respond to textual instructions.
-In this setup, the vision encoder typically remains frozen, serving solely as a feature extractor, while all learning occurs in the LLM and the projection module.
-While this paradigm effectively leverages perception to enhance reasoning, the reverse direction, i.e.
-using reasoning to improve perception, remains largely underexplored.
+A key challenge in this paradigm is the scarcity of high-quality, large-scale visual instruction data.
+Foundational models like LLaVA [103] addressed this by creating synthetic datasets: they used a powerful language model (e.g., GPT-4 [1]) to generate diverse and complex reasoning-based datasets.
+In the visual instruction tuning standard setup, the vision encoder typically remains frozen, serving solely as a feature extractor, while all learning is concentrated in the LLM and the projection module.
+While this paradigm effectively leverages perception to enhance understanding, the reverse direction, i.e.
+using understanding to improve perception, remains largely underexplored.
 Our proposed ViTP addresses this gap by inverting the traditional visual instruction tuning pipeline.
 Instead of using images to tune an LLM, we utilize an LLM to guide the pretraining of the vision encoder.
+More related work is provided in the Supplementary Material.
 3 Method
 Figure 4: A conceptual illustration of the ViTP framework.
-A ViT backbone is embedded within a large VLM and then pretrained with domain-specific instruction following objective and Visual Robustness Learning (VRL).
+A Vision Transformer (ViT) backbone is embedded within a large VLM and then pretrained with domain-specific instruction following objective and Visual Robustness Learning (VRL).
 This process instils high-level semantic understanding into the ViT.
@@
 In this section, we present our visual instruction pretraining (ViTP) framework, a novel paradigm designed to pretrain a ViT backbone.
-As illustrated in Figure 2, ViTP fundamentally differs from traditional pretraining by leveraging the reasoning capabilities of modern VLMs.
+As illustrated in Figure 2, ViTP fundamentally differs from traditional pretraining by leveraging the understanding capabilities of modern VLMs.
 The entire process is driven by a “visual instruction following objective”, where the ViT learns to extract features that help an LLM answer questions about an image.
@@
 3.1 Visual Instruction Following Objective
-The central hypothesis of ViTP is that a ViT can learn more potent and relevant features if its training is guided by high-level reasoning.
+The central hypothesis of ViTP is that a ViT can learn relevant and expressive features if its training is guided by high-level understanding.
 To achieve this, we frame the pretraining as a visual instruction-following task.
@@
 We then continue its training on our curated domain-specific datasets.
-This offers two primary advantages: 1) It leverages the VLM’s vast pre-existing knowledge of general visual and linguistic patterns, providing a robust initialization.
-2) This strong starting point significantly accelerates convergence during domain-specific pretraining, enhancing computational efficiency.
+Such a continual pretraining strategy leverages the VLM’s vast pre-existing knowledge of general visual and linguistic patterns, providing a robust initialization.
+It significantly accelerates convergence during domain-specific pretraining, enhancing computational efficiency.
 Let the raw dataset be $\mathcal{D}_{\text{raw}}=\{(I,Q,R)\}$, where $I$ is an image, $Q$ is a text query (instruction), and $R$ is the ground-truth text response.
@@
 3.1.3 LLM-based Supervision
-The combined sequence $S_{llm}$ is processed by the LLM, which acts as a reasoning engine to interpret the visual information from $x_{i}$ in the context of the instruction $x_{t}$.
+The combined sequence $S_{llm}$ is processed by the LLM, which acts as a understanding engine to interpret the visual information from $x_{i}$ in the context of the instruction $x_{t}$.
 The model then auto-regressively generates an output sequence $O$.
@@
     Task Capability Alignment: The instruction-following tasks in the pretraining data should foster abilities required downstream.
-For instance, for downstream object detection tasks, including visual grounding and fine-grained VQA during pretraining empowers the backbone with localization and spatial reasoning capabilities.
+For instance, for downstream object detection tasks, including visual grounding and fine-grained VQA during pretraining empowers the backbone with localization and spatial understanding capabilities.
   4. 4.
@@
 Once ViTP is complete, the pretrained ViT backbone is then extracted to serve as a powerful backbone network.
-For downstream applications, this pretrained ViT is integrated into a standard ViT-Adapter [18] and combined with task-specific heads for detection, segmentation, or other tasks.
+For downstream applications, the pretrained ViT is integrated into a standard ViT-Adapter [23] and combined with task-specific heads.
 The entire model is then finetuned on the target downstream dataset.
 This transfer learning approach leverages the rich, instruction-aware representations learned during ViTP, enabling faster adaptation and superior performance on specialized tasks.
-Table 1: Overview of downstream task datasets for remote sensing, including object detection (horizontal and oriented), segmentation (semantic and instance), and change detection.
-Task Dataset Modal Box Format Images Classes
-  Object Detection DIOR [63] RGB Hori. Box 23,463 20
-DIOR-R [25] RGB Ori. Box 23,463 20
-DOTA-v2.0 [146] RGB Ori. Box 11,268 18
-SARDet-100K [77] SAR Hori. Box 116,598 6
-SSDD [167] SAR Hori. Box 1,160 1
-RSAR [170] SAR Ori. Box 95,842 6
-Semantic and Instance Segmentation iSAID [143] RGB Mask 2,806 15
-LoveDA [139] RGB Mask 5,987 7
-UAVid [99] RGB Mask 5,510 8
-SSDD [167] SAR Polygons 1,160 1
-Change Detection SVCD [62] RGB Mask 16,000 2
-WHU [57] RGB Mask 11,456 2
-LEVIR-CD [11] RGB Mask 10,192 2
-S2Looking [115] RGB Mask 5,000 2
-Medical Segmentation AMOS2022 [58] CT Mask 19,310 14
-BraTS2021 [30] MRI Mask 76,467 3
-CovidQUEx [124] X-ray Mask 5,826 2
-3.5 Implementation Details
-Training a VLM from scratch conventionally requires a multi-stage pipeline: vision–language contrastive pretraining, projector alignment, and large-scale instruction tuning.
-To bypass the prohibitive cost of these initial phases, we instead bootstrap from a publicly available, high-capacity VLM.
-Specifically, we initialise our weights from InternVL-2.5 [19], which uses a custmized ViT-Large backbone [20] and a Qwen2 [2] language model.
-All pretraining are performed on 8× NVIDIA A40 (48 GB) GPUs with a global batch size of 128, and all downstream task finetunings are conducted on $8\times$ NVIDIA RTX3090 (24 GB) GPUs.
-For ViTP pretraining, we use the AdamW optimizer with a learning rate of 2e-5 and a cosine decay schedule over 8,000 training steps.
-To accelerate both pre-training and downstream finetuning, Flash-Attention [29] is integrated into every ViT self-attention layer, yielding extra throughput gain without altering convergence behaviour.
 4 Experiments
 This section offers a rigorous assessment of ViTP’s finetuning capabilities on various downstream tasks from both remote-sensing and medical imaging domains.
-Dataset statistics and task specifications are summarized in Table 1.
-This section presents our experimental setup, a thorough analysis of the results, and an ablation study.
+We present our experimental setup, a thorough analysis of the results, and an ablation study.
 We demonstrate ViTP’s superior performance compared to existing state-of-the-art methods and analyze its efficiency and robustness.
 The best score is indicated in bold, while the second-best score is underlined.
+Implementation details, dataset descriptions, additional experiments on general-domain datasets, and extended analyzes are provided in the Supplementary Material.
 4.1 Remote Sensing
-Table 2: Composition of the pretraining dataset for remote sensing.
-       Dataset Size Sample Rate Tasks
-  Mini-InternVL [38] 1394k 0.03 Caption,VQA,OCR
-RSVQA [94] 100k 0.1 VQA
-FIT_RS [98] 100k 0.1 VQA
-GeoChat [61] 64k 2 VG
-VRSBench [72] 38k 5 VG
-RSVG [123] 5.5k 10 VG
-DIOR-RSVG [162] 27k 8 VG
-ISPRS_SAR [56] 1.5k 1 CLS
-SAR_Sentinel-1&2 [131] 16k 1 CLS
-VHM [109] 223k 1 Caption,VQA,CLS
-LevirCCcaptions [87] 50k 0.5 Caption
-GAIA [160] 33k 1 Caption
-Million-AID [95] 920k 0.05 Caption,CLS
-For the ViTP pretraining phase, we leverage a diverse set of publicly available image-text paired instruction datasets, adhering to the principles outlined in Section 3.3.
-The pretraining datasets are listed in Table 2.
-In the dataset recipe, Million-AID [95], GAIA [160], LevirCCcaptions [87], VHM [109], RSVQA [94], and FIT_RS [98] are large-scale remote sensing visual instruction datasets, providing diverse remote sensing scenes and visual tasks.
-ISPRS_SAR [56] and SAR Sentinel-1&2 [131] datasets specifically address the unique challenges of Synthetic Aperture Radar (SAR) data.
-We construct question-answer pairs related to object identification, scene interpretation, and attribute recognition in SAR images using these existing SAR classification datasets.
-This ensures modality alignment and task relevance for SAR-specific downstream applications.
-GeoChat [61], DIOR-RSVG [162], RSVG [123], and VRSBench [72] datasets primarily focus on visual grounding tasks within the remote sensing domain, providing rich instructional guidance for spatial reasoning, object attribute recognition, and target localization.
-The mini-InternVL [38] dataset contributes to general vision-language understanding capabilities, ensuring that ViTP retains broad applicability and does not overfit to highly specialized remote sensing patterns.
-The comprehensive pretraining dataset recipe enables ViTP to learn robust, instruction-aware visual representations that are highly relevant to diverse remote sensing tasks.
 Table 1: Object detection performance (mAP %) on optical remote sensing datasets.
@@
 4.1.1 Object detection
-Object detection in remote sensing [146, 149, 78, 120, 177] involves identifying and precisely localizing objects of interest (e.g., vehicles, ships, and bridges) within aerial or satellite imagery.
-This task is crucial for applications such as urban planning and disaster monitoring.
-Challenges include vast scale variations of objects, arbitrary orientations, dense object distributions, and complex backgrounds.
-To evaluate the applicability of our proposed model for remote sensing object detection tasks under various scenarios, we conducted experiments on the following datasets: DIOR [63], DIOR-R [25], DOTA-v2.0 [146], SARDet-100K [77], RSAR [170], and SSDD [167].
-We use the standard COCO evaluation metrics (mean Average Precision, mAP) to evaluate the performance of the models.
-Following previous practice [41, 170, 77], we use the oriented R-CNN [149] as the default detector for oriented object detection tasks and Cascade R-CNN [8] for horizontal object detection tasks.
-For training and testing, we resized all optical datasets (DIOR, DIOR-R, and DOTA-v2) to a standard size of $\times1024$ pixels.
-For the DOTA-v2 dataset specifically, we adopted a single-scale approach, cropping images into $\times1024$ patches with a 200-pixel overlap to handle its original large size.
-The SAR datasets, SARDet-100K and RSAR, were resized to $\times800$ pixels.
 Table 2: Object detection performance (mAP %) on SAR datasets.
@@
 For DOTA-v2.0, which is particularly challenging due to a large quantity of small objects and dense scenes, ViTP achieves a new state-of-the-art of 60.23 mAP.
-This significantly surpasses previous top performers like BillionFM (58.69), highlighting ViTP’s superior ability to handle complex spatial relationships and arbitrary object orientations, likely due to its instruction-following pretraining that fosters fine-grained localization capabilities.
+This significantly surpasses previous top performers like BillionFM (58.69), highlighting ViTP’s superior ability to handle complex spatial relationships and arbitrary object orientations.
 To be noticed, SkySense [49] pretraining requires significantly more computational resources, over $17\times$ the GPU hours of ViTP, ViTP still outperforms SkySense.
@@
 These remarkable improvements on SAR datasets demonstrate ViTP’s strong generalization capabilities and its effectiveness in handling challenging modalities.
-Table 5: Semantic segmentation performance (mIoU %) on optical datasets.
+Table 3: Semantic segmentation (mIoU %) on optical datasets.
 model iSAID LoveDA model UAVid
@@
 ViTP 71.14 54.28 ViTP 73.39
-4.1.2 Semantic Segmentation
-Semantic segmentation in remote sensing [143, 141, 167] involves classifying each pixel in an image into a predefined category, such as land cover types (e.g., forest, water, urban area) or specific objects.
-This pixel-level understanding is vital for environmental monitoring, urban planning, and resource management.
-We evaluated ViTP’s performance on segmentation tasks using the following widely-used remote sensing datasets: iSAID [143], LoveDA [139], UAVid [99], and SSDD [167].
-Following previous practice [41, 122], we use the UperNet [147] as the default segmentor for semantic segmentation tasks (iSAID, LoveDA, and UAVid) and Mask-RCNN [48] for instance segmentation tasks (SSDD).
-We report the mean Intersection over Union (mIoU) as the evaluation metric for semantic segmentation and Average Precision (AP) for instance segmentation tasks.
-Table 5 and Table 6 demonstrate ViTP’s strong performance in remote sensing semantic segmentation.
-It sets new state-of-the-art performance on iSAID, LoveDA, UAVid, and SSDD datasets.
-Especially for UAVid and SSDD datasets, ViTP surpasses previous state-of-the-art by a large margin.
-The consistent improvements across these semantic segmentation benchmarks confirm that ViTP’s instruction-following pretraining on complex fine-grained regional understanding questions enables the ViT backbone to capture rich semantic information, which is crucial for accurate pixel-level classification.
 Table 4: Detailed object detection and instance segmentation performance on the SSDD Dataset (SAR modality).
@@
 ViTP 70.80 97.80 86.60 65.90 96.80 81.80
+4.1.2 Semantic Segmentation
+Table 3 and Table 4 demonstrate ViTP’s strong performance in remote sensing semantic segmentation.
+It sets new state-of-the-art performance on iSAID, LoveDA, UAVid, and SSDD datasets.
+Especially for UAVid and SSDD datasets, ViTP surpasses previous state-of-the-art by a large margin.
+The consistent improvements across these semantic segmentation benchmarks confirm that ViTP’s instruction-following pretraining on complex fine-grained regional understanding questions enables the ViT backbone to capture rich semantic information, which is crucial for accurate pixel-level classification.
 Table 5: Change detection performance (F1-Score %) on SVCD, WHU, LEVER-CD and S2Looking datasets.
@@
 RDPNet [14] 97.20 - 91.20 -
+SDACD [104] 97.34
 WNet [145] 97.56 91.25 90.67 -
@@
 4.1.3 Change Detection
-Change detection in remote sensing focuses on identifying and characterizing differences in the state of an object or phenomenon by observing it at different times.
-This bitemporal analysis is fundamental for applications such as urban sprawl monitoring, deforestation tracking, and urbanization analysis.
-The primary challenges in this task include handling variations in illumination and atmospheric conditions between image acquisitions, precise image co-registration, and distinguishing meaningful semantic changes from irrelevant ones.
-To assess ViTP’s capabilities in this domain, we conduct experiments on three widely-recognized public datasets: SVCD [62], LEVIR-CD [11] and WHU-CD [57].
-Following the common practice, we employ a simple Siamese UperNet-based framework [147] as the default detector and report the F1-Score as our primary evaluation metric.
 The comprehensive results, presented in Table 5, show that ViTP achieves state-of-the-art performance across all of SVCD, WHU, LEVIR-CD and S2Looking datasets.
@@
 4.2 Medical Imaging
-Table 8: Composition of the pretraining dataset for medical imagery.
-       Dataset Size Sample Rate Tasks
-  Mini-InternVL [38] 1394k 0.02 Caption,VQA,OCR
-GMAI-MMBench [158] 5k 20 VQA
-Open-i [31] 7k 5 Caption
-Huatuo-OA [15] 647k 0.1 Caption
-Huatuo-VQA [15] 647k 0.5 VQA
-PMC-OA [86] 1647k 0.05 Caption
-PMC-VQA [169] 227k 0.5 VQA
-OmniMedical [52] 89k 1 VQA
-Quilt-1M [54] 723k 0.1 Caption
-Quilt-Instruct [114] 147 1 VQA
-Similar to the approach for remote sensing, we collected several open-source medical image-text paired instruction datasets.
-The pretraining datasets are listed in Table 8.
-Among them, Huatuo [15], PMC [86], and OmniMedical [52] are large-scale, general medical visual instruction datasets covering various medical imaging modalities, including CT, MRI, ultrasound, and others.
-These datasets reframe various downstream tasks into a question-answering format, augmented with millions of descriptive captions, enabling the model to build comprehensive knowledge across diverse medical imaging modalities.
-Moreover, since medical image segmentation is a key downstream task, we also include the GAMI-MMBench [158] dataset.
-It provides VQA based on ROIs and segmentation masks, which aids in aligning the model with dense prediction tasks.
-The Open-i [31] and Quilt [54] datasets are included to compensate for the relative scarcity of X-ray and pathology images in other general medical collections.
-Finally, as with the remote sensing domain, mini-InternVL [38] is incorporated to enhance general vision-language understanding.
-4.2.1 Semantic Segmentation
-Semantic segmentation in medical image analysis involves classifying each pixel in an image into a predefined category, such as an organ (e.g., lung, kidney, heart) or a specific lesion.
-This pixel-level understanding is vital for clinical diagnosis, disease analysis, and patient prognosis.
-We evaluated ViTP’s performance on segmentation tasks using three challenging medical datasets: AMOS2022 [58], BraTS2021 [30], and ConvidQUEx [124].
-These datasets encompass three primary clinical modalities: CT (AMOS2022), MRI (BraTS2021), and chest X-ray (ConvidQUEx), and covering both organ and lesion segmentation scenarios.
-We use UperNet [147] as the default segmentation head.
-We report the mean Dice Score (mDice) as the evaluation metric for all segmentation tasks.
 Table 6 demonstrates the strong performance of ViTP in medical semantic segmentation.
@@
 Furthermore, we investigate the unique advantages of the ViTP paradigm over other pre-training methods during the finetuning stage, specifically focusing on its data efficiency and robustness to various data corruptions.
-Table 10: Validation of the data recipe. Performance on the RSAR Benchmark confirms the necessity of our comprehensive data curation strategy.
+Table 7: Ablation of the data recipe. Performance on RSAR confirms the necessity of our comprehensive data curation strategy.
 Pretrain Paradigm RSAR mAP
@@
 5.2 Impact of Pretraining Steps
-Figure 5: Effect of Pretraining Duration. RSAR mAP improves with more pretraining steps before saturating at approximately 8k steps.
+Figure 5: Effect of Pretraining Duration. RSAR mAP improves with more pretraining steps before saturating at $\sim$8k steps.
 The duration of pre-training is a significant factor influencing downstream task performance.
@@
 This trend indicates that more extensive instruction tuning enables the model to learn progressively better feature representations.
-However, the performance curve begins to plateau around 8k steps, with no further improvements observed thereafter.
+However, the performance curve begins to plateau around 8k steps.
 Consequently, we set the number of ViTP pre-training steps to 8k by default for our main model, balancing performance with computational cost.
@@
 5.4 Impact of Language Model Size
-This study investigates the influence of the Large Language Model’s (LLM) capacity on ViTP’s overall performance.
+This study investigates the influence of the LLM’s capacity on ViTP’s overall performance.
 As shown in Figure 7, as the LLM size increases, the downstream performance on the RSAR dataset steadily degrades.
 This phenomenon is analogous to the findings in Masked Autoencoders [55] pretrain paradigm, which highlight the importance of a lightweight decoder.
-It indicates that for the challenging SAR modality, a more powerful LLM might compensate for suboptimal visual features extracted by the Vision Transformer (ViT).
+It indicates that for the challenging SAR modality, a more powerful LLM might compensate for suboptimal visual features extracted by the ViT.
 This could reduce the optimization pressure on the ViT, preventing it from learning the most effective modality-specific representations.
@@
 SkySense [49] 400 74.27 70.91
-ViTP 23 75.08 70.54
+ViTP 23 75.08 71.14
 5.5 Training Efficiency
-A key advantage of the ViTP paradigm is its exceptional pretraining efficiency, which lowers the barrier for developing specialized foundation models.
+A key advantage of the ViTP paradigm is its pretraining efficiency, which lowers the barrier for developing specialized foundation models.
 As detailed in Table 8, ViTP pretraining completes in one day on 8 A40 GPUs.
-This is dramatically faster than competing paradigms.
+This is dramatically faster than competing paradigms and achieves better downstream performance.
 For instance, ViTP is 2.6x faster than the efficient MIM-based method (Scale-MAE) and over 17x faster than the high-performing but computationally intensive contrastive learning method, SkySense.
+6 Conclusion
+This paper addressed the underexplored, top-down pathway from high-level understanding to low-level perception in vision foundation models.
+We introduced Visual insTruction Pretraining (ViTP), a novel paradigm that leverages the high-level semantic understanding of a Vision-Language Model to directly guide the feature learning of a ViT backbone.
+Complemented by our Visual Robustness Learning (VRL), ViTP compels the vision model to learn rich, robust, and task-relevant representations from downstream-specific instructional data.
+Our extensive experiments on 16 challenging remote sensing and medical imaging benchmarks validated the effectiveness of the proposed ViTP method.
+ViTP not only establishes new SOTA performance across a diverse range of downstream tasks but also proves remarkably computationally efficient.
+We believe this work opens a promising new avenue for a deeper integration of high-level understanding into the core of visual feature learning.
+Appendix A More Related Work
+A.1 Remote sensing Perception Foundation Model
+In the remote sensing domain, pretraining strategies are tailored to address unique challenges such as large scale variations, multi-spectral imagery, and domain-specific object categories [49, 132, 82, 140].
+Supervised methods like SAMRS [157] and MSFA [92] achieve strong performance by leveraging annotated remote sensing datasets.
+To exploit the abundance of unlabeled data, MIM-based approaches such as RingMo [142], SatMAE [32], and Scale-MAE [132] are proposed, specifically designed to model the dense small objects, multi-modal or multi-scale characteristics of satellite imagery.
+Meanwhile, contrastive learning has been explored through both image-text and image-image formulations.
+Models like GeoRSCLIP [192] and RemoteCLIP [102] leverage large-scale image-text pairs to enable zero-shot classification capabilities.
+Image-based contrastive methods such as CACo [116] and Skysense [49] focus on learning discriminative features from unlabeled remote sensing imagery.
+Hybrid approaches like CMID [124] and GFM [120] further integrate the strengths of both MIM and contrastive learning to learn more comprehensive and robust representations.
+A.2 Medical Images Perception Foundation Model
+Pretraining vision foundation models is particularly crucial in medical imaging, where labeled data is scarce and expensive to obtain.
+A prominent line of research focuses on adapting the general-purpose Segment Anything Model (SAM) [73] to the medical domain.
+Adaptations range from finetuning SAM on large-scale medical datasets, as in MedSAM [115], to pretraining SAM-like architectures from scratch on medical images, as demonstrated by SAM-Med2D [29].
+Other frameworks, such as IMIS-SAM [31], aim to enhance the interactive capabilities of SAM for clinical workflows.
+Concurrently, specialized MIM pretraining methods like MedMAE [50] and S3D [155] are developed to better capture the unique characteristics of medical data.
+A common limitation across these domain-specific foundation models is their heavy reliance on tailored architectural or training designs, which restricts their transferability to other domains.
+In contrast, our proposed ViTP framework offers a more generalizable pretraining strategy that can be seamlessly adapted to various downstream domains by simply curating the corresponding visual instruction datasets.
+A.3 Continual Pretraining
+Continual pretraining was first popularized in NLP, where Gururangan et al.
+[51] demonstrated that continuing pretraining on in-domain data significantly improves model performance.
+This concept is also successfully adapted to computer vision.
+For instance, CSPT [189] first pretrains on ImageNet and then continues pretraining on a target remote sensing dataset using an MIM objective.
+TOV [146] adopts a curriculum strategy, freezing early layers pretrained on natural images while finetuning deeper layers on specialized data.
+To reduce computational overhead, RemoteCLIP [102] initializes from pretrained CLIP [129] weights and continues pretraining on domain-specific data.
+Similarly, MSFA [92] adopts a multi-stage strategy: pretraining on ImageNet, followed by training on optical remote sensing detection datasets, and finally finetuning on SAR detection data.
+Inspired by these successes, we adopt a continual pretraining approach, initializing our VLM from the well-established InternVL [24] model.
+However, our method differs from prior work in two key aspects: (1) to mitigate catastrophic forgetting of general visual knowledge, we incorporate a small proportion of general-domain data into the specialized training corpus; (2) we introduce Visual Robustness Learning as an additional regularizer to enhance the model’s robustness during continual pretraining.
+Appendix B Implementation Details
+Training a VLM from scratch conventionally requires a multi-stage pipeline: vision–language contrastive pretraining, projector alignment, and large-scale instruction tuning.
+To bypass the prohibitive cost of these initial phases, we instead bootstrap from a publicly available, high-capacity VLM.
+Specifically, we initialise our weights from InternVL-2.5 [24], which uses a custmized ViT-Large backbone [25] and a Qwen2 [3] language model.
+All pretraining are performed on 8× NVIDIA A40 (48 GB) GPUs with a global batch size of 128, and all downstream task finetunings are conducted on $8\times$ NVIDIA RTX3090 (24 GB) GPUs.
+For ViTP pretraining, we use the AdamW optimizer with a learning rate of 2e-5 and a cosine decay schedule.
+For remote sensing and medical image foundation models, we apply a schedual of 8,000 training steps, however for the general domain foundation model, we pretrain for 16,000 steps.
+To accelerate both pre-training and downstream finetuning, Flash-Attention [34] is integrated into every Vision Transformer self-attention layer, yielding extra throughput gain without altering convergence behaviour.
+Appendix C Experimental Datasets
+C.1 Pretraining Datasets
+For the ViTP pretraining phase, we leverage a diverse set of publicly available image-text paired instruction datasets, adhering to the principles outlined in Section “Pretraining Dataset Recipe”.
+C.1.1 General Domain
+The pretraining datasets are listed in Table 9.
+In the dataset recipe, ShareGPT4V [21], DVQA [69], ChartQA [118], AI2D [71], DocVQA [119], and GeoQA+ [11] are diverse visual question answering (VQA) datasets, providing rich instructional data for complex reasoning, chart and document understanding, and geospatial question answering.
+The SynthDoG-EN [72] dataset specifically addresses the unique challenges of Optical Character Recognition (OCR) tasks.
+ShareGPT4V [21] also provides visual grounding (VG) data, enhancing spatial localization capabilities.
+GAIA [180] and Million-AID [110] contribute foundational captioning (Caption) and classification (CLS) tasks, ensuring the model develops robust general-purpose vision-language understanding.
+This comprehensive pretraining dataset recipe enables the model to learn robust, instruction-aware visual representations that are highly relevant to diverse general domain tasks.
+Table 9: Composition of the pretraining dataset for general domain.
+       Dataset Size Sample Rate Tasks
+  ShareGPT4V [21] 767k 1 VQA,VG
+DVQA [69] 200k 1 VQA
+ChartQA [118] 18k 1 VQA
+AI2D [71] 12k 1 VQA
+DocVQA [119] 10k 1 VQA
+GeoQA+ [11] 72k 1 VQA
+SynthDoG-EN [72] 30k 1 OCR
+GAIA [180] 33K 0.2 Caption
+Million-AID [110] 920k 0.01 Caption,CLS
+C.1.2 Remote Sensing Domain
+The pretraining datasets are listed in Table 10.
+In the dataset recipe, Million-AID [110], GAIA [180], LevirCCcaptions [101], VHM [128], RSVQA [109], and FIT_RS [113] are large-scale remote sensing visual instruction datasets, providing diverse remote sensing scenes and visual tasks.
+ISPRS_SAR [65] and SAR Sentinel-1&2 [151] datasets specifically address the unique challenges of Synthetic Aperture Radar (SAR) data.
+We construct question-answer pairs related to object identification, scene interpretation, and attribute recognition in SAR images using these existing SAR classification datasets.
+This ensures modality alignment and task relevance for SAR-specific downstream applications.
+GeoChat [74], DIOR-RSVG [183], RSVG [143], and VRSBench [85] datasets primarily focus on visual grounding tasks within the remote sensing domain, providing rich instructional guidance for spatial understanding, object attribute recognition, and target localization.
+The mini-InternVL [46] dataset contributes to general vision-language understanding capabilities, ensuring that ViTP retains broad applicability and does not overfit to highly specialized remote sensing patterns.
+The comprehensive pretraining dataset recipe enables ViTP to learn robust, instruction-aware visual representations that are highly relevant to diverse remote sensing tasks.
+Table 10: Composition of the pretraining dataset for remote sensing domain.
+       Dataset Size Sample Rate Tasks
+  Mini-InternVL [46] 1394k 0.03 Caption,VQA,OCR
+RSVQA [109] 100k 0.1 VQA
+FIT_RS [113] 100k 0.1 VQA
+GeoChat [74] 64k 2 VG
+VRSBench [85] 38k 5 VG
+RSVG [143] 5.5k 10 VG
+DIOR-RSVG [183] 27k 8 VG
+ISPRS_SAR [65] 1.5k 1 CLS
+SAR_Sentinel-1&2 [151] 16k 1 CLS
+VHM [128] 223k 1 Caption,VQA,CLS
+LevirCCcaptions [101] 50k 0.5 Caption
+GAIA [180] 33k 1 Caption
+Million-AID [110] 920k 0.05 Caption,CLS
+Table 11: Composition of the pretraining dataset for medical imagery domain.
+       Dataset Size Sample Rate Tasks
+  Mini-InternVL [46] 1394k 0.02 Caption,VQA,OCR
+GMAI-MMBench [178] 5k 20 VQA
+Open-i [37] 7k 5 Caption
+Huatuo-OA [18] 647k 0.1 Caption
+Huatuo-VQA [18] 647k 0.5 VQA
+PMC-OA [100] 1647k 0.05 Caption
+PMC-VQA [190] 227k 0.5 VQA
+OmniMedical [60] 89k 1 VQA
+Quilt-1M [63] 723k 0.1 Caption
+Quilt-Instruct [134] 147k 1 VQA
+C.1.3 Medical Imagery Domain
+Similar to the approach for remote sensing, we collected several open-source medical image-text paired instruction datasets.
+The pretraining datasets are listed in Table 11.
+Among them, Huatuo [18], PMC [100], and OmniMedical [60] are large-scale, general medical visual instruction datasets covering various medical imaging modalities, including CT, MRI, ultrasound, and others.
+These datasets reframe various downstream tasks into a question-answering format, augmented with millions of descriptive captions, enabling the model to build comprehensive knowledge across diverse medical imaging modalities.
+Moreover, since medical image segmentation is a key downstream task, we also include the GAMI-MMBench [178] dataset.
+It provides VQA based on ROIs and segmentation masks, which aids in aligning the model with dense prediction tasks.
+The Open-i [37] and Quilt [63] datasets are included to compensate for the relative scarcity of X-ray and pathology images in other general medical collections.
+Finally, as with the remote sensing domain, mini-InternVL [46] is incorporated to enhance general vision-language understanding.
+C.2 Domain-Specific Finetuning Datasets
+Table 12: Overview of downstream task datasets.
+Task Dataset Modal Box Format Images Classes
+  General Domain
+  Detection and Segmentation COCO [99] RGB Polygons 118,287 80
+COCO [99] RGB Hori. Box 118,287 80
+ADE20K [198] RGB Mask 20,210 150
+  Remote Sensing
+  Object Detection DIOR [77] RGB Hori. Box 23,463 20
+DIOR-R [30] RGB Ori. Box 23,463 20
+DOTA-v2.0 [166] RGB Ori. Box 11,268 18
+SARDet-100K [92] SAR Hori. Box 116,598 6
+SSDD [188] SAR Hori. Box 1,160 1
+RSAR [191] SAR Ori. Box 95,842 6
+Semantic and Instance Segmentation iSAID [164] RGB Mask 2,806 15
+LoveDA [161] RGB Mask 5,987 7
+UAVid [114] RGB Mask 5,510 8
+SSDD [188] SAR Polygons 1,160 1
+Change Detection SVCD [75] RGB Mask 16,000 2
+WHU [66] RGB Mask 11,456 2
+LEVIR-CD [16] RGB Mask 10,192 2
+S2Looking [135] RGB Mask 5,000 2
+  Medical Image
+  Medical Segmentation AMOS2022 [67] CT Mask 19,310 14
+BraTS2021 [36] MRI Mask 76,467 3
+CovidQUEx [144] X-ray Mask 5,826 2
+C.2.1 Natural Image Detection and Segmentation
+Object detection and segmentation in natural scenes are fundamental tasks in computer vision, serving as benchmarks for general-purpose visual understanding.
+Object detection involves localizing and classifying multiple objects within everyday photographs, while segmentation provides pixel-level masks for precise delineation.
+These tasks are crucial for applications ranging from autonomous driving to augmented reality.
+We evaluate ViTP’s versatility on standard natural image benchmarks using two complementary datasets: COCO [99] for object detection and instance segmentation, and ADE20K [198] for semantic segmentation.
+Following established protocols [55, 127], we adopt Mask R-CNN [57] as the default framework for COCO tasks with a 12-epoch training schedule and standard data augmentations.
+For ADE20K semantic segmentation, we employ UperNet [167] with a 160k iteration training strategy.
+We report standard COCO metrics (mAP@0.5:0.95) for detection and segmentation tasks on COCO, and mean Intersection over Union (mIoU) for ADE20K semantic segmentation.
+C.2.2 Remote Sensing Object Detection
+Object detection in remote sensing [166, 169, 93, 139, 199] involves identifying and precisely localizing objects of interest (e.g., vehicles, ships, and bridges) within aerial or satellite imagery.
+This task is crucial for applications such as urban planning and disaster monitoring.
+Challenges include vast scale variations of objects, arbitrary orientations, dense object distributions, and complex backgrounds.
+To evaluate the applicability of our proposed model for remote sensing object detection tasks under various scenarios, we conducted experiments on the following datasets: DIOR [77], DIOR-R [30], DOTA-v2.0 [166], SARDet-100K [92], RSAR [191], and SSDD [188].
+We use the standard COCO evaluation metrics (mean Average Precision, mAP) to evaluate the performance of the models.
+Following previous practice [49, 191, 92], we use the oriented R-CNN [169] as the default detector for oriented object detection tasks and Cascade R-CNN [10] for horizontal object detection tasks.
+For training and testing, we resized all optical datasets (DIOR, DIOR-R, and DOTA-v2) to a standard size of $\times1024$ pixels.
+For the DOTA-v2 dataset specifically, we adopted a single-scale approach, cropping images into $\times1024$ patches with a 200-pixel overlap to handle its original large size.
+The SAR datasets, SARDet-100K and RSAR, were resized to $\times800$ pixels.
+C.2.3 Remote Sensing Semantic Segmentation
+Semantic segmentation in remote sensing [164, 162, 188] involves classifying each pixel in an image into a predefined category, such as land cover types (e.g., forest, water, urban area) or specific objects.
+This pixel-level understanding is vital for environmental monitoring, urban planning, and resource management.
+We evaluated ViTP’s performance on segmentation tasks using the following widely-used remote sensing datasets: iSAID [164], LoveDA [161], UAVid [114], and SSDD [188].
+Following previous practice [49, 142], we use the UperNet [167] as the default segmentor for semantic segmentation tasks (iSAID, LoveDA, and UAVid) and Mask-RCNN [57] for instance segmentation tasks (SSDD).
+We report the mean Intersection over Union (mIoU) as the evaluation metric for semantic segmentation and Average Precision (AP) for instance segmentation tasks.
+C.2.4 Remote Sensing Change Detection
+Change detection in remote sensing focuses on identifying and characterizing differences in the state of an object or phenomenon by observing it at different times.
+This bitemporal analysis is fundamental for applications such as urban sprawl monitoring, deforestation tracking, and urbanization analysis.
+The primary challenges in this task include handling variations in illumination and atmospheric conditions between image acquisitions, precise image co-registration, and distinguishing meaningful semantic changes from irrelevant ones.
+To assess ViTP’s capabilities in this domain, we conduct experiments on three widely-recognized public datasets: SVCD [75], LEVIR-CD [16] and WHU-CD [66].
+Following the common practice, we employ a simple Siamese UperNet-based framework [167] as the default detector and report the F1-Score as our primary evaluation metric.
+C.2.5 Medical Imaging Semantic Segmentation
+Semantic segmentation in medical image analysis involves classifying each pixel in an image into a predefined category, such as an organ (e.g., lung, kidney, heart) or a specific lesion.
+This pixel-level understanding is vital for clinical diagnosis, disease analysis, and patient prognosis.
+We evaluated ViTP’s performance on segmentation tasks using three challenging medical datasets: AMOS2022 [67], BraTS2021 [36], and ConvidQUEx [144].
+These datasets encompass three primary clinical modalities: CT (AMOS2022), MRI (BraTS2021), and chest X-ray (ConvidQUEx), and covering both organ and lesion segmentation scenarios.
+We use UperNet [167] as the default segmentation head.
+We report the mean Dice Score (mDice) as the evaluation metric for all segmentation tasks.
+Table 13: Comparison of pretraining paradigms for end-to-end finetuning.
+Results are shown for object detection ($AP_{box}$) and instance segmentation ($AP_{mask}$) on COCO val using Mask R-CNN (1x schedule), and semantic segmentation (mIoU) on ADE20K using UperNet (160k iterations).
+All models use a ViT-Large backbone.
+Pretrain Paradigm Model COCO val ADE20K (mIoU)
+(APbox) (APmask)
+  Supervised IN. Cls. [40] 49.3 43.9 49.9
+DeiT III  [152] 48.7 41.1 52.0
+Contrastive Learning CLIP [129] 51.3 - 52.2
+MoCov3 [26] 49.3 44.0 49.1
+DINOv2 [127] 53.4 46.8 55.0
+Masked Image Modelling BEiT [7] 53.3 47.1 53.3
+MAE [55] 53.3 47.2 53.6
+4M [121] 53.7 46.4 53.4
+Understanding ViTP 53.9 46.7 55.8
+Appendix D More Experiments
+D.1 General Domains
+To validate the efficacy of our proposed ViTP paradigm, we conducted comprehensive experiments against SOTA pretraining methods in two rigorous settings: end-to-end finetuning and frozen-backbone feature evaluation.
+All comparisons utilize a ViT-Large backbone for fair assessment.
+As shown in Table 13, ViTP achieves good performance, posting top results on COCO object detection (APbox 53.9) and ADE20K semantic segmentation (mIoU 55.8).
+It significantly outperforms leading methods from all major pretraining paradigms, including supervised, contrastive learning, and Masked Image Modeling.
+On COCO instance segmentation, ViTP remains highly competitive (APmask 46.7).
+This consistent superiority highlights that the features learned via our top-down approach are highly effective for downstream adaptation.
+To more directly assess the intrinsic quality of the learned visual representations, we performed semantic segmentation on ADE20K with the frozen Vision Transformer backbone.
+As detailed in Table 14, ViTP (L/16) achieves a 51.9 mIoU, surpassing other strong multimodal foundation models like AIMv2 [45] (51.5) and various SigLIP-based architectures.
+ViTP’s leading performance in both end-to-end finetuning and frozen-feature evaluation settings provides compelling evidence for the efficacy of our top-down pretraining paradigm.
+This demonstrates that the VLM’s understanding-based objective implicitly trains the Vision Transformer to learn more powerful and generalizable representations, confirming that top-down understanding significantly benefits bottom-up perception.
+Table 14: Semantic segmentation performance (mIoU) on ADE20K using an UperNet [167] architecture with a frozen ViT backbone.
+   Model ViT Size ADE20K (mIoU)
+  SigLIP [182] So/14 400M 35.2
+SigLIP2 [69] So/16 400M 35.3
+LLaVA-SigLIP [76] So/14 400M 39.9
+COMP-SigLIP [22] So/14 400M 49.5
+AIMv2 [45] L/14 300M 51.5
+COMP-AIMv2 [22] L/14 300M 51.0
+ViTP L/16 300M 51.9
 Table 15: Data efficiency on the RSAR Benchmark. ViTP demonstrates significantly stronger performance in low-data regimes compared to MIM and contrastive methods.
@@
 ViTP 73.37 70.56 69.54 67.74 70.58 69.05 68.97 70.85 67.86 67.10 67.23 66.64 71.87 69.00 4.37
-5.6 Data Efficiency
+D.2 Data Efficiency
 We investigate ViTP’s performance in low-data regimes by finetuning it on subsets of the RSAR training dataset, ranging from 2% to 20%.
@@
 For instance, with only 2% of the training data, ViTP achieves 46.98 mAP, substantially outperforming SatMAE (37.90 mAP) and RemoteCLIP (34.78 mAP).
-Notably, ViTP finetuned on just 20% of the data (67.07 mAP) achieves performance comparable to or exceeding many state-of-the-art methods trained on the full dataset (as shown in Table 4).
+Notably, ViTP finetuned on just 20% of the data (67.07 mAP) achieves performance comparable to or exceeding many state-of-the-art methods trained on the full dataset (as shown in main paper Table 2).
 This superior data efficiency suggests that the rich semantic priors acquired during instruction tuning enable the model to generalize more effectively from a limited number of examples.
-5.7 Model Robustness
+D.3 Model Robustness
 While most research datasets contain curated clean imagery, real-world remote sensing images are often degraded by atmospheric conditions (e.g., clouds, haze), sensor noise, or processing artifacts.
@@
 This confirms that our instruction tuning paradigm, enhanced with VRL, is a highly effective technique for learning robust and comprehensive feature representations.
-6 Limitation and Future Work
+Figure 8: Comparison of self-attention activation maps from the third ViT layer for a given query patch (red cross).
+The models were pretrained using ViTP (ours), CLIP, and DINOv2.
+Left: General nature images.
+Right: Abstract concept images that require a higher degree of human-like understanding.
+Figure 9: Self-attention activation maps for a given query patch (red cross) from the ViTP pretrained Vision Transformer.
+Left: ViT is pretrained with remote sensing domain data.
+Right: ViT is pretrained with medical domain data.
+Appendix E Visualization of Self-Attention Maps
+To provide a qualitative analysis of the features learned by our ViTP backbone, we visualize its internal self-attention mechanisms.
+This comparison assesses how a ViT backbone, when pretrained with our Visual Instruction Pretraining (ViTP) , interprets semantic relationships within an image compared to backbones pretrained with leading visual foundation model pretraining methods: CLIP [129] and DINOv2 [127].
+As shown in Figure 8 and Figure 9, we compute the self-attention activation maps for a specific query patch (marked with a red cross).
+These maps show which other image regions the model “attends to” when processing the information from that single query patch.
+In our analysis, we observed the ”artifacts phenomenon” (as discussed in previous works [35, 68]), where some layers produce diffuse attention.
+We found that the third layer of the Vision Transformer provides the clearest and most ideal representation of interpretable attention across all three methods, so we visualize the outputs exclusively from this layer.
+In examples like the elephants, sheep, and giraffes, the query patch (red cross) prompts the ViTP backbone to attend to other, non-local parts of the image that are semantically related.
+This capability is even more pronounced in abstract images where the ”object” is a concept.
+We also provide more visualization on domain specific images like remote sensing and medical images in Figure 9.
+This analysis confirms that our ViTP successfully imbue the ViT backbone with a high-level semantic understanding, enabling it to group visual information based on conceptual meaning rather than just low-level features.
+Appendix F Limitation and Future Work
 Despite the strong performance and efficiency of ViTP, we acknowledge several limitations that open promising avenues for future research.
@@
 Extending the ViTP framework to other data modalities, such as video and 3D point clouds, represents another exciting frontier.
-Adapting the instruction-following objective to these domains could unlock new capabilities in temporal reasoning and 3D spatial understanding.
-7 Conclusion
-In this paper, we addressed the underexplored, top-down pathway from high-level reasoning to low-level perception in vision foundation models.
-We introduced Visual insTruction Pretraining (ViTP), a novel paradigm that leverages the high-level semantic reasoning of a Vision-Language Model to directly guide the feature learning of a ViT backbone.
-Complemented by our Visual Robustness Learning (VRL), ViTP compels the vision model to learn rich, robust, and task-relevant representations from downstream-specific instructional data.
-Our extensive experiments on 16 challenging remote sensing and medical imaging benchmarks validated the effectiveness of the proposed ViTP method.
-ViTP not only established new state-of-the-art performance across a diverse range of downstream tasks but also proved to be remarkably computationally efficient.
-We believe this work opens a promising new avenue for a deeper integration of high-level reasoning into the core of visual feature learning.
+Adapting the instruction-following objective to these domains could unlock new capabilities in temporal understanding and 3D spatial understanding.